### v2

# TorchCP: A Python Library for Conformal Prediction

Abstract
Conformal Prediction (CP) has attracted great attention from the research community due to its strict theoretical guarantees. However, researchers and developers still face challenges of applicability and efficiency when applying CP algorithms to deep learning models. In this paper, we introduce TorchCP, a comprehensive PyTorch-based toolkit to strengthen the usability of CP for deep learning models. TorchCP implements a wide range of post-hoc and training methods of conformal prediction for various machine learning tasks, including classification, regression, GNN, and LLM. Moreover, we provide user-friendly interfaces and extensive evaluations to easily integrate CP algorithms into specific tasks. Our TorchCP toolkit, built entirely with PyTorch, enables high-performance GPU acceleration for deep learning models and mini-batch computation on large-scale datasets. With the LGPL license, the code is open-sourced at https://github.com/ml-stat-Sustech/TorchCP and will be continuously updated.
Keywords: Conformal Prediction, Uncertainty Quantification, Machine Learning
1 Introduction
Conformal prediction (CP) is a powerful statistical framework that provides reliable measures of uncertainty for machine learning predictions. With conformal prediction, the model yields prediction sets or intervals designed to contain the ground truths with a specified coverage probability  (Vovk et al., 2005; Shafer and Vovk, 2008; Balasubramanian et al., 2014; Angelopoulos et al., 2020; Manokhin, 2022). In recent years, conformal prediction algorithms have increasingly been integrated into deep learning models, enabling trustworthy predictions for high-stake scenarios in the real world. The trend raises new challenges of applicability and efficiency in implementing conformal prediction algorithms across various tasks of deep learning, such as time series, graph neural networks, and large language models (Zargarbashi et al., 2023; De Toni et al., 2024; Quach et al., 2024; Dey et al., 2024).
Given the increasing popularity of deep learning, it is indispensable to develop an extensive toolkit to enable easy, fast, and flexible integration of CP algorithms in various deep learning applications. There are existing libraries of uncertainty quantification that contain implementations of some CP algorithms, such as Puncc (Mendil et al., 2023), crepes(Boström, 2022), and Fortuna (Detommaso et al., 2024). However, those implementations are typically limited in the scope of application, only supporting classification and regression tasks. It is essential to develop a comprehensive CP library that supports a wider range of deep learning tasks, such as graph node classification, ordinal classification, and text generation with large language models. Besides, most of the current libraries lack support for GPU acceleration, which is crucial for deep learning models, and do not facilitate mini-batch computation for large datasets. This leads to suboptimal computational efficiency and scalability, particularly for large-scale applications. We summarize a feature comparison of current libraries in Table 1.
Table 1: Open-source libraries with support of conformal prediction.
Library Task Support
Classification Regression Graph LLM Training GPU Mini-Batch Backend
nonconformist [^1] https://github.com/donlnz/nonconformist. Numpy
TorchUQ [^2] https://github.com/TorchUQ/torchuq?tab=readme-ov-file. PyTorch
crepes (Boström, 2022) Numpy
Puncc (Mendil et al., 2023) Numpy
MAPIE (Cordier et al., 2023) Numpy
Fortuna (Detommaso et al., 2024) JAX
TorchCP (ours) PyTorch
In this paper, we present TorchCP, a Python library of conformal prediction for deep learning, built on PyTorch with strong GPU acceleration. TorchCP addresses the above limitations through the following crucial features: 1) Comprehensiveness. TorchCP implements a wide range of state-of-the-art CP algorithms, encompassing both post-hoc and training methods. In addition to conventional classification and regression tasks, the framework extends its functionality to various deep learning scenarios, including graph node classification, large language models, and ordinal classification, making it a versatile toolkit for modern machine learning applications. 2) Scalability. TorchCP is designed in a modular fashion and provides user-friendly interfaces for CP algorithms. Researchers can integrate and experiment with various CP algorithms through modular components. The framework’s plug-in architecture also enables the application of CP training algorithms to any existing loss function, offering exceptional flexibility for machine learning practitioners. 3) Efficiency. TorchCP delivers high-performance CP algorithms with strong GPU acceleration. This native PyTorch implementation enables mini-batch computation with high efficiency, providing seamless integration with deep learning models.
In this project, we open-source TorchCP, accompanied by detailed documentation and tutorials, to facilitate the easy, flexible, and efficient implementation of CP algorithms in deep learning applications. We believe that TorchCP will promote the adoption and advancement of conformal prediction within the deep learning community.
Figure 1: The scope of TorchCP.
2 The Architecture Design of TorchCP
2.1 Overview of Implementations
TorchCP implements a wide range of state-of-the-art conformal prediction algorithms, supporting both training approaches and post-hoc methods for various deep learning tasks (Figure 1). Specifically, for classification tasks, TorchCP provides various non-conformity score functions (Gammerman and Vovk, 2007; Löfström et al., 2015; Sadinle et al., 2019; Romano et al., 2020; Angelopoulos et al., 2020; Huang et al., 2024), CP algorithms (Vovk, 2012; Tibshirani et al., 2019; Ding et al., 2024), and conformal trainer  (Stutz et al., 2022; Dey et al., 2024; Xi et al., 2024; Liu et al., 2024). For regression tasks, TorchCP implements the standard conformal regression algorithm (Lei et al., 2018), multiple variants of Conformal Quantile Regression (CQR) (Romano et al., 2019; Kivaranovic et al., 2020; Sesia and Candès, 2020), and various conformal algorithms for time series problems (Gibbs and Candes, 2021; Xu and Xie, 2021).
Moreover, TorchCP provides conformal prediction algorithms for node classification of graphs and large language models (LLMs). Specifically, TorchCP implements various variants of APS that are specifically designed for graph structure (Wijegunawardana et al., 2020; Clarkson, 2023; Zargarbashi et al., 2023; Huang et al., 2023; Song et al., 2024), and the certified risk control algorithm for text generation task of LLMs (Quach et al., 2024).
Finally, to facilitate the evaluation and selection of CP methods, TorchCP offers extensive metrics for assessing the marginal/conditional coverage and size efficiency of prediction sets produced by different CP methods. This enables researchers to benchmark the superiority of novel approaches and identify the most suitable existing methods for specific tasks.
2.2 Modular Design
To improve extensibility and flexibility, TorchCP adopts a highly modularized architecture in implementing conformal prediction methods. For each task, TorchCP implements three main components of conformal prediction algorithms: Trainer, Score and Predictor. We illustrate the modular design of TorchCP on the left of Figure 2. The details of the three components are introduced below.
  • Trainer module defines those training algorithms with loss functions for improving conformal prediction, like ConfTr (Stutz et al., 2022) and C-Adapter (Liu et al., 2024).
  • Score module defines those non-conformity scores that quantify how different a new sample is from the training data, e.g., THR (Sadinle et al., 2019) ,APS (Gibbs and Candes, 2021), SAPS (Huang et al., 2024) and CQR (Romano et al., 2019).
  • Predictor module defines the conformal prediction procedure, including the functions of calibration and evaluation. For example, two common-used conformal predictors are Split CP (Vovk et al., 2005) and Class-wise CP (Vovk, 2012). A conformal predictor is defined with a score and a trained model from the trainer. The evaluate function outputs the two measures of predictors: average set size and empirical coverage rate. We provide more metrics for measuring the predictors in utils/metrics.
Example Code for Conformal Prediction. On the right side of Figure 2, we present an example code snippet for multi-class classification. First, we prepare a calibration dataset and a test dataset (Lines 5-6), along with a pre-trained PyTorch model (Lines 9-10). Next, we define a split predictor using the THR score and the trained model (Line 14). The calibration procedure is conducted by running the calibrate function with the cal_dataloader and the significance level $\alpha$ (Line 17). Finally, we obtain the evaluation results through the evaluate function (Lines 20-21). This modular design is compatible with various deep learning tasks, allowing for easy implementation of new conformal prediction algorithms.
Figure 2: An illustration of the TorchCP framework.
3 Conclusion and Future work
This paper presents TorchCP, a Python toolbox for conformal prediction research on deep learning models, built on the PyTorch Library with strong GPU acceleration. TorchCP implements a wide range of state-of-the-art conformal prediction algorithms and provides user-friendly interfaces along with comprehensive evaluation functions. Besides, we provide extensive documents, tutorials, and unit tests (100% test coverage) to improve the usability of TorchCP. The toolbox can support future research and applications of conformal prediction algorithms on deep learning models. We believe that TorchCP library will promote the adoption and advancement of conformal prediction within the deep learning community. In the future, TorchCP will be continuously updated with new algorithms, paradigms, and evaluation protocols. We also welcome any contributions for TorchCP.
Appendix A Details of conformal methods in TorchCP
A.1 Split CP
Conformal prediction (Balasubramanian et al., 2014; Manokhin, 2022; Angelopoulos et al., 2023, 2024) is a statistical framework that generates uncertainty sets/intervals containing ground-truth label with a desired probability. Split Conformal Prediction is the most widely-used version of the conformal prediction procedure. Specifically, split conformal prediction divides a given dataset into two disjoint subsets: one for training the base classifier and the other for conformal calibration. We next outline the main process of split conformal prediction:
  1. 1.
    Divide a given dataset into two disjoint subsets: a training fold $\mathcal{D}_{tr}$ and a calibration fold $\mathcal{D}_{cal}$, with $|\mathcal{D}_{cal}|$ being $n$;
  2. 2.
    Train a deep learning model on the training dataset $\mathcal{D}_{tr}$, and define a non-conformity score function $V(\bm{x},y)$;
  3. 3.
    Compute $\widehat{Q}_{1-\alpha}$ as the $\frac{\lceil(n+1)(1-\alpha)\rceil}{n}$ quantile of the calibration scores $\left\{V(\bm{x}_{i},y_{i}):(\bm{x}_{i},y_{i})\in\mathcal{D}_{cal}\right\}$, where $\widehat{Q}_{1-\alpha}$ is defined by
    $\widehat{Q}_{1-\alpha}:=\inf\{Q\in\mathbb{R}:\frac{|\{i:V(\bm{x}_{i},y_{i})\leq Q\}|}{n}\geq\frac{\lceil(n+1)(1-\alpha)\rceil}{n}\};$
  4. 4.
    Use the conformal threshold $\widehat{Q}_{1-\alpha}$ to generate a prediction set for a new instance $\bm{x}_{n+1}$:
    $\mathcal{C}(\bm{x}_{n+1})=\{y:V(\bm{x}_{n+1},y)\leq\widehat{Q}_{1-\alpha}\}.$
A.2 Classification
Here, we consider multi-class classification with $K$ classes. Let $\mathcal{X}\subset\mathbb{R}^{d}$ be the input space and $\mathcal{Y}:=\{1,\dots,K\}$ be the label space. We use $(X,Y)\sim\mathcal{P}_{\mathcal{X}\mathcal{Y}}$ to denote a random data pair satisfying a joint data distribution $\mathcal{P}_{\mathcal{X}\mathcal{Y}}$ and $f:\mathcal{X}\rightarrow\mathbb{R}^{K}$ to denote a classification neural network. Thus, the classifier $\pi:\mathcal{X}\rightarrow\Delta^{K-1}$is defined as $\sigma\circ f$, where $\Delta^{K-1}$ is a ($K$-1)-dimensional probability simplex and $\sigma$ is a normalization function such as the softmax function. Let $\pi_{y}(x)$ be the probability of $y$ conditional on $x$, i.e., $P(y|x)$. In the following, we briefly describe methods of conformal prediction supported by TorchCP.
KNN (Gammerman and Vovk, 2007): KNN calculates the score based on the k closest training examples in the feature space. The score is calculated by dividing two sums: the numerator is the sum of the k shortest distances from an input point to other points that share its same label, while the denominator is the sum of the k shortest distances from that input point to points that have different labels than it does.
Margin (Löfström et al., 2015): The margin score is constructed on a well-known concept that a prediction with a larger margin is generally considered to be more conforming. Consequently, the margin score is defined by $V{(\bm{x},y)}:=\pi_{y}-\max_{j\in\mathcal{Y}:{j}\neq y}\pi_{j}.$
THR (Sadinle et al., 2019): The THR measures the similarity between the example and the data space by the conditional probability $P(Y|X)$: $V(\bm{x},y):=1-\pi_{y}(\bm{x})$.
APS (Romano et al., 2020): The score of Adaptive Prediction Sets (APS) calculates the non-conformity score using the cumulative sum of sorted softmax probabilities:
$V(\bm{x},y,u):=\sum\limits_{y^{\prime}\in\mathcal{Y}}\pi_{y^{\prime}}(\bm{x})\mathds{1}(r_{f}(\bm{x},y^{\prime})<r_{f}(\bm{x},y))+u\cdot\pi_{y}(\bm{x}),$
where $r_{f}(\bm{x},y)$ denotes the rank of $\pi_{y}(\bm{x})$ among the descending softmax probabilities, and $u$ is an independent random variable satisfying a uniform distribution on $[0,1]$.
RAPS (Angelopoulos et al., 2020): RAPS adds a regularization to penalize noisy tail probabilities and regularize the number of samples in the uncertainty set:
$V(\bm{x},y,u;\pi):=\sum\limits_{y^{\prime}\in\mathcal{Y}}\pi_{y^{\prime}}(\bm{x})\mathds{1}(r_{f}(\bm{x},y^{\prime})<r_{f}(\bm{x},y))+u\cdot\pi_{y}(\bm{x})+\eta\cdot(r_{f}(\bm{x},y)-k_{reg})^{+},$
where $\eta$ represents the weight of regularization, $k_{reg}\geq 0$ are regularization hyper-parameters and $(z)^{+}$ denotes the positive part of $z$.
TOPK (Angelopoulos et al., 2020): TOPK is a score function based solely on the ranking information: $V(\bm{x},y,u):=\sum\limits_{y^{\prime}\in\mathcal{Y}}\mathds{1}(r_{f}(\bm{x},y^{\prime})<r_{f}(\bm{x},y))+u$.
SAPS (Huang et al., 2024): The SAPS method mitigates the issue of probability miscalibration by only retaining the highest probability value and discarding all others:
$V(\bm{x},y,u):=\left\{\begin{array}[]{ll}u\cdot\pi_{max}(\bm{x}),\qquad\text{if\quad}r_{f}(\bm{x},y)=1,\\ \pi_{max}(\bm{x})+(r_{f}(\bm{x},y)-2+u)\cdot\lambda,\quad\text{else},\end{array}\right.$
where $\lambda$ is a hyperparameter representing the weight of ranking information and $\pi_{max}(\bm{x})$ denotes the maximum softmax probability.
Class-wise CP (Vovk, 2012): Class-wise CP partitions the calibration data based on their labels $y$. The non-conformity scores’ quantiles are then computed separately for each class using the corresponding calibration subset.
Clustered CP (Ding et al., 2024): Clustered CP extends Class-wise CP by applying split conformal prediction to class clusters rather than individual classes. It achieves cluster-conditional coverage as an approximation of class-conditional guarantees.
Weighted CP (Tibshirani et al., 2019): Weighted CP addresses distribution shift in Conformal Prediction, where calibration and test data follow different distributions. The core idea is that data points closer to the test point in the feature space provide more valuable information about it.
ConfTr (Stutz et al., 2022): ConfTr simulates the CP process based on end-to-end training by splitting mini-batches into calibration and prediction sets and designing differentiable conformal prediction steps: $\mathcal{L}_{\mathrm{ConfTr}}=\max\left(0,\sum_{k=1}^{K}\mathcal{C}_{\theta,k}(\bm{x})-\kappa\right),$ where $\theta$ is the trainable parameters of deep learning models and $\kappa$ is the target set size. The definition of $\mathcal{C}_{\theta,k}(\bm{x})$ is given by $\sigma((V(\bm{x},k)-\widehat{Q}_{1-\alpha})/\tau)$ where $\tau$ is a temperature hyperparameter and $\sigma$ is the sigmoid function.
ConfTS (Xi et al., 2024): ConfTs is a training method to optimize the $T$ of temperature scaling for improved efficiency of prediction sets. The objective of ConfTS is to minimize the efficiency gap, i.e., $\widehat{Q}_{1-\alpha}-V(\bm{x},y)$. Formally, the loss function of ConfTS is:
$\mathcal{L}_{\mathrm{ConfTS}}=(\widehat{Q}_{1-\alpha}-V(\bm{x},y))^{2}$
C-Adapter (Liu et al., 2024): C-Adapter proposes to optimize an adapter with a Conformal Discriminative (CD) Loss. The CD loss is constructed to improve the $\alpha$-free efficiency of prediction sets: $\mathcal{L}_{\mathrm{CD}}=\frac{1}{K}\sum_{j\in\mathcal{Y}}\{\sigma(V(\bm{x},y)-V(\bm{x},j))\}$.
A.3 Regression
In the context of regression, the goal is to predict a continuous target variable $y\in\mathbb{R}$ based on an input $\bm{x}\in\mathcal{X}\subset\mathbb{R}^{d}$, where $(X,Y)\sim\mathcal{P}_{\mathcal{X}\mathcal{Y}}$ follows a joint data distribution $\mathcal{P}_{\mathcal{X}\mathcal{Y}}$. Let $f:\mathcal{X}\to\mathcal{F}$ denote a regression model trained on a dataset $\{(\bm{x_{i}},y_{i})\}_{i=1}^{n}$, where $\mathcal{F}$ represents the space of model outputs, $\bm{x}_{i}$ are the features and $y_{i}$ are the observed target values.
ABS (Lei et al., 2018): The ABS score $V(\bm{x},y):=|f(\bm{x})-y|$ is designed to measure the absolute deviation of a prediction $f(\bm{x})$ from the observed target value $y$, where $f(\bm{x})$ is a point estimate of given $\{\bm{x}\}$. Then the prediction set for a new instance $\bm{x}_{n+1}$ is constructed by $\mathcal{C}(\bm{x}_{n+1})=[f(\bm{x}_{n+1})-\widehat{Q}_{1-\alpha},f(\bm{x}_{n+1})+\widehat{Q}_{1-\alpha}].$
CQR (Romano et al., 2019): The CQR score $V(\bm{x},y):=\max(f(\bm{x})_{\alpha/2}-y,\;y-$$f(\bm{x})_{1-\alpha/2})$ measures the maximum absolute deviation between the true value $y$ and the predicted quantile interval $[f(\bm{x})_{\alpha/2},f(\bm{x})_{1-\alpha/2}]$. Then the prediction set for a new instance $\bm{x}_{n+1}$ is constructed by: $\mathcal{C}(\bm{x}_{n+1})=[f(\bm{x})_{\alpha/2}-\widehat{Q}_{1-\alpha},\;f(\bm{x})_{1-\alpha/2}+\widehat{Q}_{1-\alpha}].$
CQRR (Sesia and Candès, 2020): The CQRR score extends the standard CQR method by introducing a scaling mechanism. This score normalizes the deviations between the true value $y$ and the predicted quantile interval $[f(\bm{x})_{\alpha/2},f(\bm{x})_{1-\alpha/2}]$ by the width of the interval itself. Formally, for a data pair $(\bm{x},y)$, the CQRR score is defined as $V(\bm{x},y):=\max\left(\frac{f(\bm{x})_{\alpha/2}-y}{\hat{s}_{\alpha}(\bm{x})},\;\frac{y-f(\bm{x})_{1-\alpha/2}}{\hat{s}_{\alpha}(\bm{x})}\right),$ where $\hat{s}_{\alpha}(\bm{x}):=f(\bm{x})_{1-\alpha/2}-f(\bm{x})_{\alpha/2}+\epsilon$ is the scaling factor, and $\epsilon>0$ is a small constant added to prevent division by zero. The prediction set is given by $\mathcal{C}(\bm{x}_{n+1})=\left[f(\bm{x}_{n+1})_{\alpha/2}-\widehat{Q}_{1-\alpha}\cdot\hat{s}_{\alpha}(\bm{x}_{n+1}),\;f(\bm{x}_{n+1})_{1-\alpha/2}+\widehat{Q}_{1-\alpha}\cdot\hat{s}_{\alpha}(\bm{x}_{n+1})\right].$
CQRM (Sesia and Candès, 2020): The CQRM score extends the CQRR score by introducing two scaling factors to independently normalize the deviations for the lower and upper bounds of the prediction interval. This method requires the model to output predictions at three quantile levels: $[\alpha/2,1/2,1-\alpha/2]$. The two scaling factors are defined as $\hat{s}_{\alpha,\text{lo}}(\bm{x}):=f(\bm{x})_{1/2}-f(\bm{x})_{\alpha/2}+\epsilon$ and $\hat{s}_{\alpha,\text{up}}(\bm{x}):=f(\bm{x})_{1-\alpha/2}-f(\bm{x})_{1/2}+\epsilon$. The non-conformity score for a data pair $(\bm{x},y)$ and prediction interval are then defined as:
$V(\bm{x},y):=\max\left(\frac{f(\bm{x})_{\alpha/2}-y}{\hat{s}_{\alpha,\text{lo}}(\bm{x})},\;\frac{y-f(\bm{x})_{1-\alpha/2}}{\hat{s}_{\alpha,\text{up}}(\bm{x})}\right).$
Then, the prediction set is given by
$\mathcal{C}(\bm{x}_{n+1})=\left[f(\bm{x}_{n+1})_{\alpha/2}-\widehat{Q}_{1-\alpha}\cdot\hat{s}_{\alpha,\text{lo}}(\bm{x}_{n+1}),\;f(\bm{x}_{n+1})_{1-\alpha/2}+\widehat{Q}_{1-\alpha}\cdot\hat{s}_{\alpha,\text{up}}(\bm{x}_{n+1})\right].$
CQRFM (Kivaranovic et al., 2020): The CQRFM score is similar to CQRM, as it also requires the model to output predicted quantiles at $[\alpha/2,1/2,1-\alpha/2]$. However, it normalizes the deviations from the median $f(\bm{x})_{1/2}$ by the fractional distance to the nearest quantile bounds. The two fractional scaling factors are defined in the same way as in CQRM. The non-conformity score and prediction interval are defined as:
$V(\bm{x},y):=\max\left(\frac{f(\bm{x})_{1/2}-y}{\hat{s}_{\alpha,\text{lo}}(\bm{x})},\;\frac{y-f(\bm{x})_{1/2}}{\hat{s}_{\alpha,\text{up}}(\bm{x})}\right).$
Then, the prediction set is given by
$\mathcal{C}(\bm{x}_{n+1})=\left[f(\bm{x}_{n+1})_{1/2}-\widehat{Q}_{1-\alpha}\cdot\hat{s}_{\alpha,\text{lo}}(\bm{x}_{n+1}),\;f(\bm{x}_{n+1})_{1/2}+\widehat{Q}_{1-\alpha}\cdot\hat{s}_{\alpha,\text{up}}(\bm{x}_{n+1})\right].$
R2CCP (Guha et al., 2023): R2CCP converts a regression problem into a classification problem by dividing the range space into $K$ bins. The model predicts the probability that each sample belongs to one of these $K$ bins, similar to a standard classification task. To train the regression-to-classification model, R2CCP uses a custom loss function:
$\mathcal{L}(\theta)=\sum_{i=1}^{N}\left(\sum_{k=1}^{K}\left(|y_{i}-m_{k}|^{p}\cdot\pi_{ik}\right)-\tau\cdot\sum_{k=1}^{K}\pi_{ik}\log(\pi_{ik})\right),$
where $\pi_{ik}$ is the predicted probability for the $k$-th bin, $m_{k}$ are the bin midpoints, $\tau$ is a regularization parameter, and $p$ is the exponent applied to the absolute difference between the target value and the bin midpoint. Both $\tau$ and $p$ are hyperparameters. After training the regression-to-classification model, R2CCP defines the non-conformity score as the linear interpolation of the model’s softmax probabilities. Based on these scores, R2CCP outputs the interval where all non-conformity scores are smaller than $\widehat{Q}_{1-\alpha}$.
Ensemble CP (Xu and Xie, 2021): Ensemble CP trains multiple models based on randomly sampled subsets of the data and aggregates their nonconformity scores using a specific aggregation function, such as mean or median. It can update $\widehat{Q}_{1-\alpha}$ over time, making it adaptable to dynamic time series.
ACI (Gibbs and Candes, 2021): ACI dynamically adjusts the confidence level $\alpha_{t}=\alpha_{t-1}+\gamma\left(\alpha-\text{err}_{t}\right)$ based on the historical error rates to adapt to distribution shifts, where $\alpha_{t}$ is the updated confidence level depending on time $t$, $\alpha$ is the target coverage, and $\text{err}_{t}:=\sum_{i=1}^{n}w_{i}\cdot I(y_{i}\notin\hat{C}_{i})$ is the weighted error rate at time step $t$. This dynamic updating enables ACI to maintain reliable prediction sets despite evolving data distributions.
A.4 Graph
Here, we consider node classification tasks. The Graph is represented as $\mathcal{G}=(\mathcal{V},\mathcal{E})$, where $\mathcal{V}:=\{v_{i}\}_{i=1}^{N}$ denotes the node set and $\mathcal{E}$ denotes the edge set with $|\mathcal{E}|=E$. Let $\bm{X}:=[\bm{x}_{1},\cdot\cdot\cdot,\bm{x}_{N}]^{T}$ be the node feature matrix, where $\bm{x}_{i}\in\mathbb{R}^{d}$ is a $d$-dimensional feature vector for node $v_{i}$. The label of node $v_{i}$ is $y_{i}\in\mathcal{Y}$, where $\mathcal{Y}:=\{1,2,...,K\}$ denotes the label space.
DAPS (Zargarbashi et al., 2023): DAPS proposes a diffusion-based method that incorporates neighborhood information by leveraging the network homophily. Specifically, the score function is defined as: $V(\bm{x}_{i},y,\lambda):=(1-\lambda)\hat{V}(\bm{x}_{i},y)+\frac{\lambda}{|\mathcal{N}_{i}|}\sum_{v_{j}\in\mathcal{N}_{i}}\hat{V}(\bm{x}_{j},y),$ where $\hat{V}(\cdot,\cdot)$ is the basic non-conformity score function, e.g., APS and RAPS, and $\lambda$ is a diffusion parameter.
SNAPS (Song et al., 2024): SNAPS, which is essentially an enhanced version of the DAPS, aggregates non-conformity scores of nodes with high feature similarity to ego node and one-hop graph structural neighbors. Formally, the score function is shown as: $V(\bm{x}_{i},y,\lambda,\mu):=(1-\lambda-\mu)\hat{V}(\bm{x}_{i},y)+\frac{\lambda}{\bm{D}_{s}(i,i)}\sum_{j=1}^{M}\bm{A}_{s}(i,j)\hat{V}(\bm{x}_{j},y)+\frac{\mu}{|\mathcal{N}_{i}|}\sum_{v_{j}\in\mathcal{N}_{i}}\hat{V}(\bm{x}_{j},y),$ where $\bm{A}_{s}$ is the adjacency matrix of $k$-NN graph, which measures the similarity between each node pair, and $\bm{D}_{s}$ is the degree matrix of $\bm{A}_{s}$. $M$ denotes the number of randomly selected nodes as targets used to construct $k$-NN graph. Both $\lambda$ and $\mu$ are hyperparameters, which are used to measure the importance of three parts of non-conformity scores.
NAPS (Clarkson, 2023): NAPS, which adapts the weighted variant of CP from (Barber et al., 2023), assigns a weight to adjacent nodes, and a weight of zero otherwise. NAPS then applies nonexchangeable conformal prediction with the APS scoring function in the following equation: $\mathcal{C}(\bm{x}_{n+1})=\{y\in\mathcal{Y}:V(\bm{x}_{n+1},y)\leq Q_{1-\alpha}\left(\sum_{i}^{n}w_{i}\cdot\delta_{s_{i}}+w_{n+1}\cdot\delta_{+\infty}\right)\},$ where $Q_{\tau}(\cdot)$ denotes the $\tau$-quantile of a distribution and $s_{i}=V(\bm{x}_{i},y_{i})$. Moreover, $w_{i}$ denotes a weight of $s_{i}$ and $\delta_{s_{i}}$ denotes a point mass at $s_{i}$.
CF-GNN (Huang et al., 2023): CF-GNN introduces a topology-aware output correction model, akin to GCN, which employs a conformal-aware inefficiency loss to refine predictions and improve the efficiency of post-hoc CP. For the loss, CF-GNN adopts the optimizing efficiency version of ConfTr.
A.5 LLMs
Conformal LLM (Quach et al., 2024): This paper proposes a novel conformal procedure based on risk control framework (Bates et al., 2021) for language models by reformulating the generation task as a sampling problem rather than enumerating all possible outputs. The core algorithm samples candidate outputs until a calibrated stopping rule is met, while using a rejection rule to filter out low-quality samples. The method guarantees that the output set contains at least one acceptable answer with high probability while maintaining the statistical guarantees of traditional conformal prediction.
Appendix B Benchmark Results
B.1 Classification
In Figure 3, we present the experimental results of split CP on the ImageNet dataset, using THR, APS, RAPS, and SAPS. In Figure 4, we present a comparative analysis of different conformal prediction methods on the ImageNet dataset, including THR, APS, RAPS, Split CP, Class-wise CP, and Clustered CP. Clustered CP achieves the smallest average size and best class-conditional coverage.
(a) Coverage
(b) Size
Figure 3: Results of image classification with $\alpha=0.1$ on ImageNet dataset.
(a) Coverage
(b) Size
(c) CovGap
Figure 4: Results of image classification with $\alpha=0.1$ on ImageNet dataset .
B.2 Regression
In Figure 5, we present an example of experimental results on the Community dataset (Redmond, 2002), using ABS, CQR, CQRR, CQRM, CQRFM, R2CCP with Split CP.
(a) Coverage
(b) Size
Figure 5: Results of regression with $\alpha=0.1$ on Community dataset.
B.3 Graph Node Classification
In Figure 6, we present an example of experimental results on the CoraML dataset (McCallum et al., 2000), using APS, DAPS, CF-GNN, and SNAPS. The experiments are conducted in a transductive learning setting for node classification tasks.
(a) Coverage
(b) Size
Figure 6: Results of Graph node classification with $\alpha=0.1$ on CoraML dataset.
B.4 LLM
In Table 2, we present an example prediction set, generated by conformal LLM, of the TriviaQA dataset (Joshi et al., 2017). The results show that the generation process of the prediction set stops after sample 5, because the sum of scores collectively exceeds the conformal threshold.
Table 2: Example prediction set of conformal LLM on the TriviaQA dataset.
Question Men Against the Sea and Pitcairn’s Island were two sequels to what famous novel?
Answer Score Label
Sample 1 Mutiny On The Bounty 0.5347 True
Sample 2 Treasure Island 0.5087 False
Sample 3 Treasure Island 0.5087 False
Sample 4 Mutiny on the Bounty 0.1889 True
Sample 5 Robinson Crusoe 0.1384 False
Appendix C Comparison of Computational Efficiency
We evaluated the computational efficiency of TorchCP by comparing its inference time with other existing libraries of conformal prediction. In particular, we focus on the time of calibration and inference on MNIST (LeCun et al., 2010) and ImageNet (Russakovsky et al., 2015) throughout the conformal prediction process. We leave out some libraries that cannot be implemented on the dataset without making significant changes. The results in Figure 7 show that TorchCP is the most efficient implementation among these libraries. TorchCP exhibits superior computational efficiency, attributed to its seamless integration with PyTorch’s architecture and mini-batch parallel processing capabilities.
(a) MNIST
(b) ImageNet
Figure 7: Time comparison of conformal prediction using various open-sourced libraries on MNIST and ImageNet datasets.
Appendix D Extensibility
Each module in TorchCP can be easily extended. In the following, we provide basic guidelines for customizing your own scores, predictors, trainers, and metrics.
D.1 Add new scores
Adding new scores involves two steps:
  1. 1.
    Implementing a New Score Class: new Scores should be implemented in the score module, inheriting from the BaseScore class. Using classification as an example, one could alternatively choose to override the functions: __init__(), __call__(), _calculate_all_label(), and _calculate_single_label().
  2. 2.
    Adding an Interface: After customizing the score class, register it in score/__init__.py.
D.2 Add new predictors
Similar to adding new scores, the addition of new predictors also consists of two steps:
  1. 1.
    Implementing a New Predictor Class: new Predictors should be implemented in the predictor module, inheriting from the BasePredictor class. Using classification as an examples, one could alternatively choose to override the functions: __init__(), calibrate() and predict().
  2. 2.
    Adding an Interface: After customizing the predictor class, register it in predictor/__init__.py.
D.3 Add new trainers
Adding new trainers is similar to steps of D.1 and D.2:
  1. 1.
    Implementing a New Trainer Class: new Trainers should be implemented in the trainer module, inheriting from the BaseTrainer class. Using classification as an example, one could alternatively choose to override the functions: __init__(), train() and validate(). One could pass different loss functions in the __init__() method.
  2. 2.
    Adding an Interface: After customizing the trainer class, register it in trainer/__init__.py.
D.4 Add new metrics
Adding new metrics involves two steps:
  1. 1.
    Implementing a New Metric Function: new Metrics should be implemented in the utils/metrics.py.
  2. 2.
    Adding an Interface: After customizing the metric function, register it by Registry("METRICS") in the utils/metrics.py.

### v3

@@
 Abstract
-Conformal Prediction (CP) has attracted great attention from the research community due to its strict theoretical guarantees.
-However, researchers and developers still face challenges of applicability and efficiency when applying CP algorithms to deep learning models.
-In this paper, we introduce TorchCP, a comprehensive PyTorch-based toolkit to strengthen the usability of CP for deep learning models.
-TorchCP implements a wide range of post-hoc and training methods of conformal prediction for various machine learning tasks, including classification, regression, GNN, and LLM.
-Moreover, we provide user-friendly interfaces and extensive evaluations to easily integrate CP algorithms into specific tasks.
-Our TorchCP toolkit, built entirely with PyTorch, enables high-performance GPU acceleration for deep learning models and mini-batch computation on large-scale datasets.
-With the LGPL license, the code is open-sourced at https://github.com/ml-stat-Sustech/TorchCP and will be continuously updated.
-Keywords: Conformal Prediction, Uncertainty Quantification, Machine Learning
+Conformal prediction (CP) is a robust statistical framework that generates prediction intervals or sets with guaranteed coverage probability, addressing the challenge of quantifying predictive uncertainty in deep learning.
+Despite advancements in deep learning architectures and datasets, reliable uncertainty estimation remains elusive, making CP increasingly vital.
+This paper introduces TorchCP, a PyTorch-native library designed to integrate state-of-the-art CP algorithms into deep learning tasks, including classification, regression, graph neural networks, and large language models.
+TorchCP offers a comprehensive suite of advanced methodologies, a modular design for easy customization, and full GPU-accelerated scalability.
+Released under the LGPL-3.0 license, TorchCP has gained widespread adoption with over 12,582 PyPi downloads.
+It is supported by approximately 16,132 lines of code, 564 unit tests achieving 100% coverage, and comprehensive documentation.
+By bridging statistics and computer science, TorchCP empowers researchers and practitioners to advance conformal prediction in diverse deep learning applications.
+Keywords: Conformal Prediction, Deep Learning
 1 Introduction
-Conformal prediction (CP) is a powerful statistical framework that provides reliable measures of uncertainty for machine learning predictions.
-With conformal prediction, the model yields prediction sets or intervals designed to contain the ground truths with a specified coverage probability  (Vovk et al., 2005; Shafer and Vovk, 2008; Balasubramanian et al., 2014; Angelopoulos et al., 2020; Manokhin, 2022).
-In recent years, conformal prediction algorithms have increasingly been integrated into deep learning models, enabling trustworthy predictions for high-stake scenarios in the real world.
-The trend raises new challenges of applicability and efficiency in implementing conformal prediction algorithms across various tasks of deep learning, such as time series, graph neural networks, and large language models (Zargarbashi et al., 2023; De Toni et al., 2024; Quach et al., 2024; Dey et al., 2024).
-Given the increasing popularity of deep learning, it is indispensable to develop an extensive toolkit to enable easy, fast, and flexible integration of CP algorithms in various deep learning applications.
-There are existing libraries of uncertainty quantification that contain implementations of some CP algorithms, such as Puncc (Mendil et al., 2023), crepes(Boström, 2022), and Fortuna (Detommaso et al., 2024).
-However, those implementations are typically limited in the scope of application, only supporting classification and regression tasks.
-It is essential to develop a comprehensive CP library that supports a wider range of deep learning tasks, such as graph node classification, ordinal classification, and text generation with large language models.
-Besides, most of the current libraries lack support for GPU acceleration, which is crucial for deep learning models, and do not facilitate mini-batch computation for large datasets.
-This leads to suboptimal computational efficiency and scalability, particularly for large-scale applications.
-We summarize a feature comparison of current libraries in Table 1.
-Table 1: Open-source libraries with support of conformal prediction.
-Library Task Support
-Classification Regression Graph LLM Training GPU Mini-Batch Backend
-nonconformist [^1] https://github.com/donlnz/nonconformist. Numpy
-TorchUQ [^2] https://github.com/TorchUQ/torchuq?tab=readme-ov-file. PyTorch
+Conformal prediction (CP) is a powerful statistical framework generating prediction intervals or sets that contain true labels with a specified coverage probability (Vovk et al., 2005).
+In recent years, deep learning has evolved dramatically with advanced model architectures and vast datasets, yet it struggles to reliably measure predictive uncertainty, making conformal prediction an increasingly attractive solution.
+Beyond classification and regression, conformal prediction methods have been tailored for diverse deep learning tasks, such as graph neural networks (GNNs) for molecular modeling and large language models (LLMs) for natural language understanding.
+This underscores the need for a comprehensive and efficient CP toolbox tailored for deep learning models, facilitating robust and accessible conformal prediction methods across diverse deep learning applications.
+This work presents TorchCP, a PyTorch-native library of conformal prediction for deep learning.
+Firstly, TorchCP implements a wide range of state-of-the-art CP algorithms encompassing both training and inference methods designed for deep learning tasks, e.g., classification (ordinal classification), regression (time series), GNN, and LLM.
+Secondly, the library’s modular design allows developers to easily use or customize each component (e.g., losses or scores) of the library.
+Lastly, TorchCP is fully built on PyTorch with GPU support and batch processing, enabling seamless scalability for large-scale deep learning tasks.
+The TorchCP library, released under the LGPL-3.0 license, can be installed via PyPi with the command ‘‘pip install torchcp’’ and has been downloaded 12,582 times  [^1] https://pepy.tech/projects/torchcp.
+The implementation consists of approximately 16,132 lines of code and is thoroughly tested with 564 unit tests, achieving 100% test coverage.
+We also provide comprehensive online documentation  [^2] https://torchcp.readthedocs.io/en/latest/ to guide users through TorchCP’s features.
+We hope the library can bridge the gap between statistics and computer science, enabling researchers and practitioners from both backgrounds to advance conformal prediction in deep learning.
+2 The library design and implementation
+Figure 1: An illustration of the modular design of TorchCP.
+2.1 Modular design
+For low coupling, we separate the TorchCP library into separate modules that can be easily maintained, tested, and integrated.
+For each task, the code implementation of the library contains 3 modules: Trainer, Score, and Predictor.
+Specifically, Trainer defines training algorithms for improving conformal prediction, along with loss functions.
+Score defines non-conformity score functions that quantify how different a new sample is from the training data, e.g., LAC (Sadinle et al., 2019).
+Predictor defines CP workflows (e.g., calibration and prediction), such as Split CP  [^3] A detailed introduction of Split CP is presented in Appendix A.1 (Papadopoulos et al., 2002).
+The modular design allows a straightforward extension to new tasks or new CP methods that arise in this active research field.
+We visualize the modular design of TorchCP in Figure 1, highlighting the relationships among the three modules.
+The Trainer module is used to train or fine-tune deep models, employed by the Predictor module for calibration and prediction.
+The Score module implements non-conformity score functions, enabling their flexible use across different trainers and predictors.
+This modular structure empowers users to independently select or customize trainers, score functions, and predictors for diverse conformal prediction tasks, significantly improving flexibility and maintainability.
+Additionally, TorchCP achieves strong extensibility by providing standardized module interfaces, making it straightforward to define new methods.
+Appendix B provides guidelines for adding customized modules using the provided interfaces, while Appendix E discusses the advantages of TorchCP ’s modular design.
+2.2 Implementation
+Table 1: The CP algorithms implemented in TorchCP.
+Task Component Methods
+Classification Trainer&Loss ConfTr (Stutz et al., 2022), ConfTS (Xi et al., 2025), C-Adapter (Liu et al., 2024), UA Loss (Einbinder et al., 2022), SCPO (Bellotti, 2021)
+Score Function LAC (Sadinle et al., 2019), APS (Romano et al., 2020), RAPS (Angelopoulos et al., 2020), SAPS (Huang et al., 2024), KNN (Gammerman and Vovk, 2007), Margin (Löfström et al., 2015), TOPK (Angelopoulos et al., 2020)
+Predictor Split CP (Papadopoulos et al., 2002) Class-conditional CP (Vovk, 2012), Clustered CP (Ding et al., 2024), Weighted CP (Tibshirani et al., 2019), RC3P (Shi et al., 2024)
+Regression Score Function ABS (Papadopoulos et al., 2002), NORABS (Vovk et al., 2005), CQR (Romano et al., 2019), CQRR (Sesia and Candès, 2020), CQRM (Sesia and Candès, 2020), CQRFM (Kivaranovic et al., 2020), R2CCP (Guha et al., 2023)
+Predictor Ensemble CP (Xu and Xie, 2021), ACI (Gibbs and Candes, 2021), AgACI (Zaffran et al., 2022)
+GNN Trainer&Loss CF-GNN (Huang et al., 2023)
+Score Function DAPS (Zargarbashi et al., 2023), SNAPS (Song et al., 2024)
+Predictor NAPS (Clarkson, 2023)
+LLM Predictor Conformal LLM (Quach et al., 2024)
+Methodological scope
+Table 1 outlines the CP algorithms included in TorchCP across 4 deep learning tasks: classification, regression (time series), GNN, and LLM , along with their corresponding evaluation metrics in Table 4.
+In total, TorchCP provides 6 training algorithms, 16 score functions, and 10 predictors tailored to various tasks.
+A brief overview of all supported methods and their advantages is available in Appendix A.
+Batch Processing and GPU Support
+TorchCP is a PyTorch-native library equipped with GPU acceleration and batch processing capabilities.
+It enables efficient computation on large datasets and scales seamlessly to real-world deep learning applications.
+Appendix D.2 presents empirical evidence of its efficiency advantages on large-scale datasets.
+Tutorials
+TorchCP aims to provide a user-friendly framework for conformal prediction in deep learning.
+The library contains illustrative code showcasing functionality for common tasks (See the “examples” folder  [^4] https://github.com/ml-stat-Sustech/TorchCP/tree/master/examples), and Appendix C offers a practical guide to help users identify relevant examples.
+In Appendix D, we provide reproduced results of classical CP methods, confirming the correctness of our implementation.
+Below, we present a generic example  [^5] The full code of the example is available at https://github.com/ml-stat-Sustech/TorchCP/blob/master/examples/classification_conftr_cifar100.py to show the usability of TorchCP.
+First, we employ ConfTr (Stutz et al., 2022) to train a model for conformal prediction (optional, using standard trained models also work):
+1from torchcp.classification.trainer import ConfTrTrainer
+2trainer = ConfTrTrainer(init_model, alpha=0.1) # Create a trainer with a NN
+3trained_model = trainer.train(train_loader, num_epochs=10) # Training 10 epochs
+With the trained model, users can proceed with conformal prediction by selecting an appropriate score function and predictor.
+As an example, we implement Split CP with a miscoverage rate of $\alpha=0.1$ using the LAC score :
+1predictor = SplitPredictor(score_function=LAC(), model=trained_model, alpha=0.1)
+2predictor.calibrate(cal_loader) # Compute the threshold on calibrate data
+3predict_set = predictor.predict(test_instance) # Predict the set for an instance
+3 Related library
+Conformal Prediction (CP) is a distribution-free framework for quantifying the uncertainty of machine learning model predictions.
+While some uncertainty quantification libraries, such as PUNCC (Mendil et al., 2023), implement certain CP algorithms, their limited efficiency, scalability, and algorithmic coverage restrict their use in deep learning.
+For instance, most libraries support only a narrow range of tasks, such as regression and classification, and lack support for complex applications like generative modeling or time-series forecasting.
+Additionally, the absence of GPU acceleration and batch processing in these libraries makes them inefficient for large-scale scenarios.
+In contrast, TorchCP offers a clean interface, enhanced extensibility, high efficiency, and a comprehensive set of CP algorithms.
+For example, TorchCP reduces inference time by up to 90% compared to PUNCC on large-scale datasets (see Appendix D.2 for details).
+We summarize the comparison of different libraries in Table 2 and discuss the advantages of TorchCP ’s modular design in Appendix E.
+Table 2: Feature comparison of open-source libraries supporting conformal prediction. “Training” refers to CP-specific training algorithms, and “Online” denotes online CP.
+Library Task Support Functionality Support Backend
+Classification Regression GNN LLM Training Online GPU Batch
+nonconformist (Linusson et al., 2020) Numpy
+TorchUQ (Zhao et al., 2024) PyTorch
 Crepes (Boström, 2022) Numpy
@@
 TorchCP (ours) PyTorch
-In this paper, we present TorchCP, a Python library of conformal prediction for deep learning, built on PyTorch with strong GPU acceleration.
-TorchCP addresses the above limitations through the following crucial features: 1) Comprehensiveness.
-TorchCP implements a wide range of state-of-the-art CP algorithms, encompassing both post-hoc and training methods.
-In addition to conventional classification and regression tasks, the framework extends its functionality to various deep learning scenarios, including graph node classification, large language models, and ordinal classification, making it a versatile toolkit for modern machine learning applications.
-2) Scalability.
-TorchCP is designed in a modular fashion and provides user-friendly interfaces for CP algorithms.
-Researchers can integrate and experiment with various CP algorithms through modular components.
-The framework’s plug-in architecture also enables the application of CP training algorithms to any existing loss function, offering exceptional flexibility for machine learning practitioners.
-3) Efficiency.
-TorchCP delivers high-performance CP algorithms with strong GPU acceleration.
-This native PyTorch implementation enables mini-batch computation with high efficiency, providing seamless integration with deep learning models.
-In this project, we open-source TorchCP, accompanied by detailed documentation and tutorials, to facilitate the easy, flexible, and efficient implementation of CP algorithms in deep learning applications.
-We believe that TorchCP will promote the adoption and advancement of conformal prediction within the deep learning community.
-Figure 1: The scope of TorchCP.
-2 The Architecture Design of TorchCP
-2.1 Overview of Implementations
-TorchCP implements a wide range of state-of-the-art conformal prediction algorithms, supporting both training approaches and post-hoc methods for various deep learning tasks (Figure 1).
-Specifically, for classification tasks, TorchCP provides various non-conformity score functions (Gammerman and Vovk, 2007; Löfström et al., 2015; Sadinle et al., 2019; Romano et al., 2020; Angelopoulos et al., 2020; Huang et al., 2024), CP algorithms (Vovk, 2012; Tibshirani et al., 2019; Ding et al., 2024), and conformal trainer  (Stutz et al., 2022; Dey et al., 2024; Xi et al., 2024; Liu et al., 2024).
-For regression tasks, TorchCP implements the standard conformal regression algorithm (Lei et al., 2018), multiple variants of Conformal Quantile Regression (CQR) (Romano et al., 2019; Kivaranovic et al., 2020; Sesia and Candès, 2020), and various conformal algorithms for time series problems (Gibbs and Candes, 2021; Xu and Xie, 2021).
-Moreover, TorchCP provides conformal prediction algorithms for node classification of graphs and large language models (LLMs).
-Specifically, TorchCP implements various variants of APS that are specifically designed for graph structure (Wijegunawardana et al., 2020; Clarkson, 2023; Zargarbashi et al., 2023; Huang et al., 2023; Song et al., 2024), and the certified risk control algorithm for text generation task of LLMs (Quach et al., 2024).
-Finally, to facilitate the evaluation and selection of CP methods, TorchCP offers extensive metrics for assessing the marginal/conditional coverage and size efficiency of prediction sets produced by different CP methods.
-This enables researchers to benchmark the superiority of novel approaches and identify the most suitable existing methods for specific tasks.
-2.2 Modular Design
-To improve extensibility and flexibility, TorchCP adopts a highly modularized architecture in implementing conformal prediction methods.
-For each task, TorchCP implements three main components of conformal prediction algorithms: Trainer, Score and Predictor.
-We illustrate the modular design of TorchCP on the left of Figure 2.
-The details of the three components are introduced below.
-  • Trainer module defines those training algorithms with loss functions for improving conformal prediction, like ConfTr (Stutz et al., 2022) and C-Adapter (Liu et al., 2024).
-  • Score module defines those non-conformity scores that quantify how different a new sample is from the training data, e.g., THR (Sadinle et al., 2019) ,APS (Gibbs and Candes, 2021), SAPS (Huang et al., 2024) and CQR (Romano et al., 2019).
-  • Predictor module defines the conformal prediction procedure, including the functions of calibration and evaluation.
-For example, two common-used conformal predictors are Split CP (Vovk et al., 2005) and Class-wise CP (Vovk, 2012).
-A conformal predictor is defined with a score and a trained model from the trainer.
-The evaluate function outputs the two measures of predictors: average set size and empirical coverage rate.
-We provide more metrics for measuring the predictors in utils/metrics.
-Example Code for Conformal Prediction.
-On the right side of Figure 2, we present an example code snippet for multi-class classification.
-First, we prepare a calibration dataset and a test dataset (Lines 5-6), along with a pre-trained PyTorch model (Lines 9-10).
-Next, we define a split predictor using the THR score and the trained model (Line 14).
-The calibration procedure is conducted by running the calibrate function with the cal_dataloader and the significance level $\alpha$ (Line 17).
-Finally, we obtain the evaluation results through the evaluate function (Lines 20-21).
-This modular design is compatible with various deep learning tasks, allowing for easy implementation of new conformal prediction algorithms.
-Figure 2: An illustration of the TorchCP framework.
 4 Conclusion and Future work
-This paper presents TorchCP, a Python toolbox for conformal prediction research on deep learning models, built on the PyTorch Library with strong GPU acceleration.
-TorchCP implements a wide range of state-of-the-art conformal prediction algorithms and provides user-friendly interfaces along with comprehensive evaluation functions.
-Besides, we provide extensive documents, tutorials, and unit tests (100% test coverage) to improve the usability of TorchCP.
-The toolbox can support future research and applications of conformal prediction algorithms on deep learning models.
-We believe that TorchCP library will promote the adoption and advancement of conformal prediction within the deep learning community.
-In the future, TorchCP will be continuously updated with new algorithms, paradigms, and evaluation protocols.
-We also welcome any contributions for TorchCP.
-Appendix A Details of conformal methods in TorchCP
+In this work, we introduced TorchCP, a PyTorch-native library that seamlessly integrates conformal prediction (CP) into deep learning workflows.
+Supporting a wide range of state-of-the-art CP algorithms, TorchCP leverages PyTorch’s GPU acceleration and batch processing, offering a modular design for easy customization and user-friendly documentation for accessibility.
+Looking ahead, we plan to expand TorchCP by incorporating emerging CP algorithms for evolving architectures like diffusion models and multimodal models, while optimizing for real-time performance.
+Through open-source collaboration on GitHub, we aim to foster community-driven innovation, empowering reliable AI in fields such as healthcare, finance, and autonomous systems.
+Appendix A Details of Conformal Prediction Methods in TorchCP
+In this section, we provide brief introductions to the methods implemented in TorchCP.
+Furthermore, Table 3 offers practical guidelines to help users select appropriate methods based on their specific application needs.
+Table 3: Advantages or use cases of CP methods across different tasks.
+Component Method Advantage/Use case
+Classification
+Score LAC To achieve the smallest prediction sets
+APS To approximate conditional coverage
+RAPS For smaller, more stable sets than APS
+TOPK A simple baseline with only logit ranking
+SAPS Improving efficiency, conveying instance-wise uncertainty
+Margin A non-conformity function based on prediction margin
+KNN A score function based solely on feature-space nearest neighbors without a predictive model.
+Predictor Class-conditional CP To achieve coverage for every class
+Clustered CP For class-conditional coverage with limited data per class
+Weighted CP Conformal prediction under covariate shift
+Loss & Trainer ConfTr Efficient prediction sets with conformal-aware loss
+C-Adapter Optimizing efficiency based on $\alpha$-free loss
+ConfTS Improving efficiency via temperature scaling
+Regression
+Score ABS Simple and interpretable residual-based scores
+NORABS Improving robustness of prediction intervals via noise-aware normalization
+CQR Flexible interval width via quantile regression
+R2CCP Shorter prediction intervals than CQR via recasting regression as classification
+Predictor ACI Reliable coverage under distributional shifts
+AgACI Aggregates ACI predictions with adaptive weighting
+Ensemble CP Stable prediction intervals without data splitting via model ensembling
+GNN
+Score DAPS Improves efficiency of APS using structural neighbors
+SNAPS Extends DAPS with feature neighbors for better efficiency
+NAPS Constructing prediction sets for inductive node classification via Weighted CP
+Loss & Trainer CF-GNN Reducing set size by attaching an adapter with a topology-aware inefficiency loss
+LLM
+Predictor Conformal LLM To achieve statistically guaranteed prediction sets for language models through calibrated sampling
 A.1 Split CP
-Conformal prediction (Balasubramanian et al., 2014; Manokhin, 2022; Angelopoulos et al., 2023, 2024) is a statistical framework that generates uncertainty sets/intervals containing ground-truth label with a desired probability.
-Split Conformal Prediction is the most widely-used version of the conformal prediction procedure.
-Specifically, split conformal prediction divides a given dataset into two disjoint subsets: one for training the base classifier and the other for conformal calibration.
+Conformal prediction (Papadopoulos et al., 2002; Balasubramanian et al., 2014; Manokhin, 2022; Angelopoulos et al., 2023, 2024) is a statistical framework that generates uncertainty sets/intervals containing ground-truth labels with a desired probability.
+Split Conformal Prediction (Papadopoulos et al., 2002) (Split CP) is the most widely-used version of the conformal prediction procedure.
+Specifically, Split CP divides a given dataset into two disjoint subsets: one for training the base model and the other for conformal calibration.
 We next outline the main process of split conformal prediction:
@@
   2. 2.
-    Train a deep learning model on the training dataset $\mathcal{D}_{tr}$, and define a non-conformity score function $V(\bm{x},y)$;
+    Train a model on the training dataset $\mathcal{D}_{tr}$, and then define a non-conformity score function $V(\bm{x},y)$;
   3. 3.
@@
 A.2 Classification
-Here, we consider multi-class classification with $K$ classes.
+Here, we provide a brief overview of conformal prediction methods for multi-class classification tasks as supported by TorchCP.
 Let $\mathcal{X}\subset\mathbb{R}^{d}$ be the input space and $\mathcal{Y}:=\{1,\dots,K\}$ be the label space.
@@
 Let $\pi_{y}(x)$ be the probability of $y$ conditional on $x$, i.e., $P(y|x)$.
-In the following, we briefly describe methods of conformal prediction supported by TorchCP.
-KNN (Gammerman and Vovk, 2007): KNN calculates the score based on the k closest training examples in the feature space.
-The score is calculated by dividing two sums: the numerator is the sum of the k shortest distances from an input point to other points that share its same label, while the denominator is the sum of the k shortest distances from that input point to points that have different labels than it does.
-Margin (Löfström et al., 2015): The margin score is constructed on a well-known concept that a prediction with a larger margin is generally considered to be more conforming.
-Consequently, the margin score is defined by $V{(\bm{x},y)}:=\pi_{y}-\max_{j\in\mathcal{Y}:{j}\neq y}\pi_{j}.$
-THR (Sadinle et al., 2019): The THR measures the similarity between the example and the data space by the conditional probability $P(Y|X)$: $V(\bm{x},y):=1-\pi_{y}(\bm{x})$.
+LAC (Sadinle et al., 2019): The LAC measures the similarity between the example and the data space by the conditional probability $P(Y|X)$: $V(\bm{x},y):=1-\pi_{y}(\bm{x})$.
 APS (Romano et al., 2020): The score of Adaptive Prediction Sets (APS) calculates the non-conformity score using the cumulative sum of sorted softmax probabilities:
@@
 where $\lambda$ is a hyperparameter representing the weight of ranking information and $\pi_{max}(\bm{x})$ denotes the maximum softmax probability.
-Class-wise CP (Vovk, 2012): Class-wise CP partitions the calibration data based on their labels $y$.
+KNN (Gammerman and Vovk, 2007): KNN calculates the score based on the k closest training examples in the feature space.
+The score is calculated by dividing two sums: the numerator is the sum of the k shortest distances from an input point to other points that share its same label, while the denominator is the sum of the k shortest distances from that input point to points that have different labels than it does.
+Margin (Löfström et al., 2015): The margin score is constructed on a well-known concept that a prediction with a larger margin is generally considered to be more conforming.
+Consequently, the margin score is defined by $V{(\bm{x},y)}:=\pi_{y}-\max_{j\in\mathcal{Y}:{j}\neq y}\pi_{j}.$
+Class-conditional CP (Vovk, 2012): Class-conditional CP partitions the calibration data based on their labels $y$.
 The non-conformity scores’ quantiles are then computed separately for each class using the corresponding calibration subset.
-Clustered CP (Ding et al., 2024): Clustered CP extends Class-wise CP by applying split conformal prediction to class clusters rather than individual classes.
+Clustered CP (Ding et al., 2024): Clustered CP extends Class-conditional CP by applying split conformal prediction to class clusters rather than individual classes.
 It achieves cluster-conditional coverage as an approximation of class-conditional guarantees.
@@
 The core idea is that data points closer to the test point in the feature space provide more valuable information about it.
+RC3P (Shi et al., 2024): RC3P addresses the inefficiency of class-conditional conformal prediction, which often results in large prediction sets, especially for imbalanced classification tasks.
+The core idea is to selectively perform class-wise thresholding only for labels with low top-k error, by integrating a label rank calibration step into standard class-wise conformal calibration.
 ConfTr (Stutz et al., 2022): ConfTr simulates the CP process based on end-to-end training by splitting mini-batches into calibration and prediction sets and designing differentiable conformal prediction steps: $\mathcal{L}_{\mathrm{ConfTr}}=\max\left(0,\sum_{k=1}^{K}\mathcal{C}_{\theta,k}(\bm{x})-\kappa\right),$ where $\theta$ is the trainable parameters of deep learning models and $\kappa$ is the target set size.
-The definition of $\mathcal{C}_{\theta,k}(\bm{x})$ is given by $\sigma((V(\bm{x},k)-\widehat{Q}_{1-\alpha})/\tau)$ where $\tau$ is a temperature hyperparameter and $\sigma$ is the sigmoid function.
-ConfTS (Xi et al., 2024): ConfTs is a training method to optimize the $T$ of temperature scaling for improved efficiency of prediction sets.
+The definition of $\mathcal{C}_{\theta,y}(\bm{x})$ is given by $\sigma((\widehat{Q}_{1-\alpha}-V(\bm{x},y))/\tau)$ where $\tau$ is a temperature hyperparameter and $\sigma$ is the sigmoid function.
+ConfTS (Xi et al., 2025): ConfTs is a training method to optimize the $T$ of temperature scaling for improved efficiency of prediction sets.
 The objective of ConfTS is to minimize the efficiency gap, i.e., $\widehat{Q}_{1-\alpha}-V(\bm{x},y)$.
@@
 The CD loss is constructed to improve the $\alpha$-free efficiency of prediction sets: $\mathcal{L}_{\mathrm{CD}}=\frac{1}{K}\sum_{j\in\mathcal{Y}}\{\sigma(V(\bm{x},y)-V(\bm{x},j))\}$.
+SCPO (Bellotti, 2021): SCPO proposes an approach to train the deep learning model directly by optimizing for maximum predictive efficiency.
+To enable end-to-end training, the conformal predictor is approximated by a differentiable surrogate objective.
+Gradient descent is then used to minimize the following loss:
+$\mathcal{L}_{\mathrm{SCPO}}=\sum_{k=1}^{K}\mathcal{C}_{\theta,k}(\bm{x})+\lambda\left(\mathcal{C}_{\theta,y}(\bm{x})-(1-\alpha)\right)^{2},$
+where $\lambda>0$ controls the relative importance of the two objective components.
+Uncertainty-aware (UA) Loss (Einbinder et al., 2022): Uncertainty-aware Loss introduces a novel training objective that integrates conformal calibration directly into model training by encouraging the conformity scores to follow a uniform distribution.
+The novel uncertainty-aware component of this loss is:
+$\mathcal{L}_{u}=\sup_{w\in[0,1]}\left|\mathbb{E}[\mathds{1}(V(\bm{x},y,u)-w]\right|,$
+where $V(\bm{x},y,u):=\sum\limits_{y^{\prime}\in\mathcal{Y}}\pi_{y^{\prime}}(\bm{x})\mathds{1}(r_{f}(\bm{x},y^{\prime})<r_{f}(\bm{x},y))+u\cdot\pi_{y}(\bm{x})$.
 A.3 Regression
@@
 Let $f:\mathcal{X}\to\mathcal{F}$ denote a regression model trained on a dataset $\{(\bm{x_{i}},y_{i})\}_{i=1}^{n}$, where $\mathcal{F}$ represents the space of model outputs, $\bm{x}_{i}$ are the features and $y_{i}$ are the observed target values.
-ABS (Lei et al., 2018): The ABS score $V(\bm{x},y):=|f(\bm{x})-y|$ is designed to measure the absolute deviation of a prediction $f(\bm{x})$ from the observed target value $y$, where $f(\bm{x})$ is a point estimate of given $\{\bm{x}\}$.
+ABS (Papadopoulos et al., 2002): The ABS score $V(\bm{x},y):=|f(\bm{x})-y|$ is designed to measure the absolute deviation of a prediction $f(\bm{x})$ from the observed target value $y$, where $f(\bm{x})$ is a point estimate of given $\{\bm{x}\}$.
 Then the prediction set for a new instance $\bm{x}_{n+1}$ is constructed by $\mathcal{C}(\bm{x}_{n+1})=[f(\bm{x}_{n+1})-\widehat{Q}_{1-\alpha},f(\bm{x}_{n+1})+\widehat{Q}_{1-\alpha}].$
+NORABS (Vovk et al., 2005): The NORABS score $V(\bm{x},y):=\frac{|f(\bm{x})-y|}{\hat{\sigma}(\bm{x})}$ is a normalized variant of ABS.
+It uses models $f(\bm{x})$ and $\hat{\sigma}(\bm{x})$ trained to maximize the likelihood of the data with respect to $\mathbb{E}[Y_{\text{test}}|X_{\text{test}}=\bm{x}]$ and $\sqrt{\text{Var}[Y_{\text{test}}|X_{\text{test}}=\bm{x}]}$, respectively.
+Then the prediction set for a new instance $\bm{x}_{n+1}$ is constructed by $\mathcal{C}(\bm{x}_{n+1})=[f(\bm{x}_{n+1})-\widehat{Q}_{1-\alpha}\cdot\hat{\sigma}(\bm{x}_{n+1}),f(\bm{x}_{n+1})+\widehat{Q}_{1-\alpha}\cdot\hat{\sigma}(\bm{x}_{n+1})].$
 CQR (Romano et al., 2019): The CQR score $V(\bm{x},y):=\max(f(\bm{x})_{\alpha/2}-y,\;y-$$f(\bm{x})_{1-\alpha/2})$ measures the maximum absolute deviation between the true value $y$ and the predicted quantile interval $[f(\bm{x})_{\alpha/2},f(\bm{x})_{1-\alpha/2}]$.
@@
 This dynamic updating enables ACI to maintain reliable prediction sets despite evolving data distributions.
-A.4 Graph
-Here, we consider node classification tasks.
-The Graph is represented as $\mathcal{G}=(\mathcal{V},\mathcal{E})$, where $\mathcal{V}:=\{v_{i}\}_{i=1}^{N}$ denotes the node set and $\mathcal{E}$ denotes the edge set with $|\mathcal{E}|=E$.
+AgACI (Zaffran et al., 2022): AgACI is a parameter-free conformal prediction method for time series that builds on ACI by using online expert aggregation.
+It runs multiple models in parallel with different learning rates $\gamma_{k}$ and adaptively aggregates their prediction intervals using weighted averages, where weights are based on the historical pinball loss.
+This approach eliminates the need to tune $\gamma$ manually and ensures valid coverage while improving interval efficiency, especially under temporal dependencies.
+A.4 GNN
+We consider the prediction tasks on graph data.
+The Graph is defined as $\mathcal{G}=(\mathcal{V},\mathcal{E})$, where $\mathcal{V}:=\{v_{i}\}_{i=1}^{N}$ denotes the node set and $\mathcal{E}$ denotes the edge set with $|\mathcal{E}|=E$.
 Let $\bm{X}:=[\bm{x}_{1},\cdot\cdot\cdot,\bm{x}_{N}]^{T}$ be the node feature matrix, where $\bm{x}_{i}\in\mathbb{R}^{d}$ is a $d$-dimensional feature vector for node $v_{i}$.
-The label of node $v_{i}$ is $y_{i}\in\mathcal{Y}$, where $\mathcal{Y}:=\{1,2,...,K\}$ denotes the label space.
+The label of $v_{i}$ is $y_{i}\in\mathcal{Y}$, where $\mathcal{Y}:=\{1,2,...,K\}$ denotes the label space.
 DAPS (Zargarbashi et al., 2023): DAPS proposes a diffusion-based method that incorporates neighborhood information by leveraging the network homophily.
-Specifically, the score function is defined as: $V(\bm{x}_{i},y,\lambda):=(1-\lambda)\hat{V}(\bm{x}_{i},y)+\frac{\lambda}{|\mathcal{N}_{i}|}\sum_{v_{j}\in\mathcal{N}_{i}}\hat{V}(\bm{x}_{j},y),$ where $\hat{V}(\cdot,\cdot)$ is the basic non-conformity score function, e.g., APS and RAPS, and $\lambda$ is a diffusion parameter.
+Specifically, the score function is defined as: $V(\bm{x}_{i},y,\lambda):=(1-\lambda)\hat{V}(\bm{x}_{i},y)+\frac{\lambda}{|\mathcal{N}_{i}|}\sum_{v_{j}\in\mathcal{N}_{i}}\hat{V}(\bm{x}_{j},y),$ where $\hat{V}(\cdot,\cdot)$ is the basic non-conformity score function, e.g., APS, and $\lambda$ is a diffusion parameter.
 SNAPS (Song et al., 2024): SNAPS, which is essentially an enhanced version of the DAPS, aggregates non-conformity scores of nodes with high feature similarity to ego node and one-hop graph structural neighbors.
@@
 Moreover, $w_{i}$ denotes a weight of $s_{i}$ and $\delta_{s_{i}}$ denotes a point mass at $s_{i}$.
-CF-GNN (Huang et al., 2023): CF-GNN introduces a topology-aware output correction model, akin to GCN, which employs a conformal-aware inefficiency loss to refine predictions and improve the efficiency of post-hoc CP.
+CF-GNN (Huang et al., 2023): CF-GNN introduces a topology-aware output correction model, akin to Graph Convolutional Network, which employs a conformal-aware inefficiency loss to refine predictions and improve the efficiency of post-hoc CP.
 For the loss, CF-GNN adopts the optimizing efficiency version of ConfTr.
-A.5 LLMs
-Conformal LLM (Quach et al., 2024): This paper proposes a novel conformal procedure based on risk control framework (Bates et al., 2021) for language models by reformulating the generation task as a sampling problem rather than enumerating all possible outputs.
+A.5 LLM
+Conformal LLM (Quach et al., 2024): This paper proposes a novel conformal procedure based on a risk control framework (Bates et al., 2021) for language models by reformulating the generation task as a sampling problem rather than enumerating all possible outputs.
 The core algorithm samples candidate outputs until a calibrated stopping rule is met, while using a rejection rule to filter out low-quality samples.
 The method guarantees that the output set contains at least one acceptable answer with high probability while maintaining the statistical guarantees of traditional conformal prediction.
-Appendix B Benchmark Results
-B.1 Classification
-In Figure 3, we present the experimental results of split CP on the ImageNet dataset, using THR, APS, RAPS, and SAPS.
-In Figure 4, we present a comparative analysis of different conformal prediction methods on the ImageNet dataset, including THR, APS, RAPS, Split CP, Class-wise CP, and Clustered CP.
-Clustered CP achieves the smallest average size and best class-conditional coverage.
+A.6 Metrics
+In this section, we provide a brief overview of the evaluation metrics supported by TorchCP for various predictive tasks, as summarized in Table 4.
+These metrics facilitate rigorous and fine-grained assessment of prediction sets, offering deep insights into model uncertainty.
+Table 4: Implemented evaluation metrics in TorchCP.
+Task Metrics
+Classification Coverage, Size, CovGap (Ding et al., 2024), WSC (Romano et al., 2020), SSCV (Angelopoulos et al., 2020), Violated classes (Kasa and Taylor, 2023), DiffViolation (Angelopoulos et al., 2020), , etc.
+Regression Coverage, Size
+GNN Coverage, Size, Singleton ratio (Zargarbashi et al., 2023)
+LLM SetLoss, SSCL (Quach et al., 2024)
+Appendix B Extensibility
+Each module in TorchCP can be easily extended. In the following, we provide basic guidelines for customizing your own scores, predictors, trainers, and metrics.
+B.1 Add new scores
+Adding new scores involves two steps:
+  1. 1.
+    Implementing a New Score Class: new Scores should be implemented in the score module, inheriting from the BaseScore class.
+Using classification as an example, one could alternatively choose to override the functions: __init__(), __call__(), _calculate_all_label(), and _calculate_single_label().
+  2. 2.
+    Adding an Interface: After customizing the score class, register it in score/__init__.py.
+B.2 Add new predictors
+Similar to adding new scores, the addition of new predictors also consists of two steps:
+  1. 1.
+    Implementing a New Predictor Class: new Predictors should be implemented in the predictor module, inheriting from the BasePredictor class.
+Using classification as an example, one could alternatively choose to override the functions: __init__(), calibrate(), and predict().
+  2. 2.
+    Adding an Interface: After customizing the predictor class, register it in predictor/__init__.py.
+B.3 Add new trainers
+Adding new trainers is similar to steps of B.1 and B.2:
+  1. 1.
+    Implementing a New Trainer Class: new Trainers should be implemented in the trainer module, inheriting from the BaseTrainer class.
+Using classification as an example, one could alternatively choose to override the functions: __init__(), train() and validate().
+One could pass different loss functions in the __init__() method.
+  2. 2.
+    Adding an Interface: After customizing the trainer class, register it in trainer/__init__.py.
+B.4 Add new metrics
+Adding new metrics involves two steps:
+  1. 1.
+    Implementing a New Metric Function: New metrics should be implemented in the utils/metrics.py.
+  2. 2.
+    Adding an Interface: After customizing the metric function, register it by Registry(‘‘METRICS’’) in the utils/metrics.py.
+Appendix C Example Codes for TorchCP
+TorchCP provides a collection of example codes in the “examples" folder  [^6] https://github.com/ml-stat-Sustech/TorchCP/tree/master/examples of the official codebase, designed to help users quickly get started with various conformal prediction tasks.
+In this section, we briefly introduce representative examples for the four supported tasks.
+C.1 Example codes for offline conformal prediction
+This part provides brief descriptions of example codes under the offline setting:
+  1. 1.
+    Classification.
+This example code [^7] https://github.com/ml-stat-Sustech/TorchCP/blob/master/examples/classification_splitcp_cifar100.py applies Split CP with the LAC score on CIFAR-100 using a pretrained ResNet-20.
+It evaluates the coverage and average set size of the resulting prediction sets.
+  2. 2.
+    Regression.
+The example code of regression [^8] https://github.com/ml-stat-Sustech/TorchCP/blob/master/examples/regression_cqr_synthetic.py shows the complete workflow of applying Split CP with the Conformalized Quantile Regression score (Romano et al., 2019) on a synthetic dataset in regression task, including model training, calibration, prediction interval generation, and evaluation.
+  3. 3.
+    GNN.
+The node classification example [^9] https://github.com/ml-stat-Sustech/TorchCP/blob/master/examples/gnn_transductive_coraml.py implements a Graph Convolutional Network (GCN) (Kipf and Welling, 2017) for node classification on the CoraML dataset (McCallum et al., 2000) under a transductive setting.
+Conformal prediction is then applied to generate prediction sets with guaranteed coverage.
+  4. 4.
+    LLM.
+The LLM example [^10] https://github.com/ml-stat-Sustech/TorchCP/blob/master/examples/llm_ConformalLM_TriviaQA.py demonstrates how to deploy conformal prediction in large language model generation tasks, i.e., TriviaQA (Joshi et al., 2017).
+C.2 Example codes for online conformal prediction
+Online conformal prediction means that after predicting a test point, its true label is immediately revealed, and the test sample is incorporated into the calibration set.
+TorchCP supports this setting as well, with the following examples:
+  1. 1.
+    Regression.
+In TorchCP, the online setup of the regression task is integrated into ACIPredictor (Gibbs and Candes, 2021).
+Therefore, we added an example code  [^11] https://github.com/ml-stat-Sustech/TorchCP/blob/master/examples/timeseries_aci_synthetic.py for the regression task in TorchCP.
+  2. 2.
+    Classification.
+For classification tasks, this online calibration approach can be naturally implemented in TorchCP by iteratively invoking the “calibrate()” function with new incoming data.
+To clarify this, we have added an online conformal prediction example 121212 https://github.com/ml-stat-Sustech/TorchCP/blob/master/examples/classification_splitcp_cifar100_online.py for a classification task in TorchCP.
+Appendix D Empirical Results
+This section presents empirical results of TorchCP.
+We first reproduce several classical conformal prediction methods across various tasks.
+Then, we benchmark the computational efficiency of TorchCP against several representative conformal prediction libraries.
+D.1 Evaluation results
+Classification.
+In Figure 2, we present the experimental results of Split CP on ImageNet, using LAC, APS, RAPS (penalty=0.1, kreg =0), and SAPS (weight=0.2).
+In Figure 3, we present a comparative analysis of different conformal prediction methods on the ImageNet dataset, including Split CP, Class-conditional CP, and Clustered CP with APS.
+The number of trials is 1000.
 (a) Coverage
@@
 (c) CovGap
-Figure 4: Results of image classification with $\alpha=0.1$ on ImageNet dataset .
-B.2 Regression
-In Figure 5, we present an example of experimental results on the Community dataset (Redmond, 2002), using ABS, CQR, CQRR, CQRM, CQRFM, R2CCP with Split CP.
+Figure 3: Conditional coverage results on ImageNet image classification with $\alpha=0.1$.
+Regression.
+In Figure 4, we present an example of experimental results on the Community dataset (Redmond, 2002), using ABS, CQR, CQRR, CQRM, CQRFM, R2CCP with Split CP. The number of trials is 1000.
 (a) Coverage
@@
 Figure 4: Results of regression with $\alpha=0.1$ on Community dataset.
-B.3 Graph Node Classification
-In Figure 6, we present an example of experimental results on the CoraML dataset (McCallum et al., 2000), using APS, DAPS, CF-GNN, and SNAPS.
-The experiments are conducted in a transductive learning setting for node classification tasks.
+GNN.
+In Figure 5, we present experimental results on the CoraML dataset (McCallum et al., 2000) for node classification under a transductive learning setting.
+We evaluate four methods: APS, DAPS, CF-GNN, and SNAPS, each conducted over 1000 independent trials.
+For DAPS, the diffusion parameter is set to 0.5.
+In CF-GNN, the topology-aware correction model is implemented using a two-layer GCN.
+For SNAPS, the hyperparameters are set as follows: $\lambda=\frac{1}{3}$, $\mu=\frac{1}{3}$, and $k=20$.
+These hyperparameter choices follow the default or recommended configurations reported in the respective original papers.
 (a) Coverage
@@
 Figure 5: Results of Graph node classification with $\alpha=0.1$ on CoraML dataset.
-B.4 LLM
+LLM.
 In Table 5, we present an example prediction set, generated by conformal LLM, of the TriviaQA dataset (Joshi et al., 2017).
-The results show that the generation process of the prediction set stops after sample 5, because the sum of scores collectively exceeds the conformal threshold.
-Table 2: Example prediction set of conformal LLM on the TriviaQA dataset.
+The generation process terminates after the fifth sample, as the cumulative score surpasses the conformal threshold.
+Table 5: An Example of prediction sets from Conformal LLM on TriviaQA.
 Question Men Against the Sea and Pitcairn’s Island were two sequels to what famous novel?
 Answer Score Label
-Sample 1 Mutiny On The Bounty 0.5347 True
-Sample 2 Treasure Island 0.5087 False
-Sample 3 Treasure Island 0.5087 False
-Sample 4 Mutiny on the Bounty 0.1889 True
-Sample 5 Robinson Crusoe 0.1384 False
-Appendix C Comparison of Computational Efficiency
-We evaluated the computational efficiency of TorchCP by comparing its inference time with other existing libraries of conformal prediction.
-In particular, we focus on the time of calibration and inference on MNIST (LeCun et al., 2010) and ImageNet (Russakovsky et al., 2015) throughout the conformal prediction process.
-We leave out some libraries that cannot be implemented on the dataset without making significant changes.
-The results in Figure 7 show that TorchCP is the most efficient implementation among these libraries.
-TorchCP exhibits superior computational efficiency, attributed to its seamless integration with PyTorch’s architecture and mini-batch parallel processing capabilities.
+Answer 1 Mutiny On The Bounty 0.5347 True
+Answer 2 Treasure Island 0.5087 False
+Answer 3 Treasure Island 0.5087 False
+Answer 4 Mutiny on the Bounty 0.1889 True
+Answer 5 Robinson Crusoe 0.1384 False
+D.2 Comparison of computational efficiency
+D.2.1 Classification Task
+We evaluated the computational efficiency of TorchCP in image classification tasks by comparing its calibration and inference time with several existing CP libraries.
+Specifically, we use two standard benchmarks: MNIST (LeCun et al., 2010) and ImageNet (Russakovsky et al., 2015).
+Throughout all experiments, we used Split CP with APS as the conformal prediction method.
+Moreover, TorchCP supports two data processing modes: full-batch (processing the entire dataset at once), and batch (processing data in mini-batches).
+In our experiments, we set the batch size to 128 for batch processing.
+We use NVIDIA H100 for GPU computation and AMD EPYC 9654 for CPU computation.
+MNIST.
+We compared the computational efficiency of five libraries: PUNCC, MAPIE, Fortuna, Crepes, and TorchCP (GPU-based full-batch processing).
+In the dataset, we use 48,000 samples for training, 6,000 for calibration, and 6,000 for testing.
+For TorchCP, we use a LeNet-5 model for training and inference using PyTorch on a GPU.
+For PUNCC, MAPIE, and Crepes, we obtain model predictions using PyTorch on a GPU, and then convert output probabilities to NumPy arrays for conformal calibration and inference using CPU.
+For Fortuna, we train the LeNet-5 model using the built-in training algorithm and perform model predictions with JAX on a GPU.
+We only report the total runtime for Fortuna as it ties the conformal calibration and inference processes together.
+Figure 6(a) presents the running times of the five libraries on the MINST dataset.
+The results show that TorchCP achieves the lowest calibration and inference time among the five libraries.
+Specifically, TorchCP completes calibration and inference for the entire dataset in just 0.07 seconds, using only 3.6% of the runtime needed by PUNCC (1.93 seconds).
+Moreover, the Crepes library exhibits the longest inference time on the MNIST dataset, likely due to its computation of p-values for each test sample’s non-conformity scores by comparing them against the entire calibration set.
+This substantial improvement highlights the efficiency advantage of TorchCP on small-scale datasets.
+ImageNet.
+We compared the computational efficiency of four libraries: PUNCC, MAPIE, Crepes, and TorchCP (GPU-based batch processing).
+In the dataset, we use a pre-trained ResNet-50 model from torchvision and select 25,000 images from the validation set for calibration and another 25,000 for testing.
+For TorchCP, we leverage GPU-based batch processing to efficiently run the conformal pipeline.
+For PUNCC, MAPIE, and Crepes, we conducted experiments on the CPU, as they lack batch processing support, and their GPU-based full-batch processing causes out-of-memory errors.
+Figure 6(b) shows the results of running times of three libraries on the ImageNet dataset.
+The results demonstrate that TorchCP achieves the shortest calibration and inference times, cutting approximately 90% of the time required by PUNCC (104s vs.
+995s).
+This significant enhancement underscores the scalability of TorchCP on large-scale datasets and models.
 (a) MNIST
 (b) ImageNet
-Figure 7: Time comparison of conformal prediction using various open-sourced libraries on MNIST and ImageNet datasets.
-Appendix D Extensibility
-Each module in TorchCP can be easily extended. In the following, we provide basic guidelines for customizing your own scores, predictors, trainers, and metrics.
-D.1 Add new scores
-Adding new scores involves two steps:
-  1. 1.
-    Implementing a New Score Class: new Scores should be implemented in the score module, inheriting from the BaseScore class.
-Using classification as an example, one could alternatively choose to override the functions: __init__(), __call__(), _calculate_all_label(), and _calculate_single_label().
-  2. 2.
-    Adding an Interface: After customizing the score class, register it in score/__init__.py.
-D.2 Add new predictors
-Similar to adding new scores, the addition of new predictors also consists of two steps:
-  1. 1.
-    Implementing a New Predictor Class: new Predictors should be implemented in the predictor module, inheriting from the BasePredictor class.
-Using classification as an examples, one could alternatively choose to override the functions: __init__(), calibrate() and predict().
-  2. 2.
-    Adding an Interface: After customizing the predictor class, register it in predictor/__init__.py.
-D.3 Add new trainers
-Adding new trainers is similar to steps of D.1 and D.2:
-  1. 1.
-    Implementing a New Trainer Class: new Trainers should be implemented in the trainer module, inheriting from the BaseTrainer class.
-Using classification as an example, one could alternatively choose to override the functions: __init__(), train() and validate().
-One could pass different loss functions in the __init__() method.
-  2. 2.
-    Adding an Interface: After customizing the trainer class, register it in trainer/__init__.py.
-D.4 Add new metrics
-Adding new metrics involves two steps:
-  1. 1.
-    Implementing a New Metric Function: new Metrics should be implemented in the utils/metrics.py.
-  2. 2.
-    Adding an Interface: After customizing the metric function, register it by Registry("METRICS") in the utils/metrics.py.
+Figure 6: Runtime comparison of different conformal prediction libraries.
+Choices of Batch processing.
+To assess the efficiency of two processing modes (Full-batch and batch), we conduct experiments with TorchCP on both MNIST and ImageNet, using the same data splits and models as in prior experiments.
+We utilize GPU-accelerated computing for all experiments except the full-batch mode on ImageNet, where the dataset’s size exceeds GPU memory capacity.
+Figures 7(a) and  7(b) present the running times of different data processing modes on MNIST and ImageNet, respectively.
+On the small-scale dataset - MNIST, using full-batch processing on GPU achieves the fastest implementation among the three choices, requiring only 26% computational time of that with CPU.
+When it comes to the ImageNet dataset, GPU-based batch processing completes calibration and inference in approximately 104.29 seconds, compared to 1022.55 seconds for CPU-based processing, achieving a 9.8× speedup.
+These findings suggest that GPU-based full-batch processing is optimal for small-scale datasets where GPU memory is sufficient to load the whole dataset.
+Conversely, GPU-based batch processing is preferable for large-scale datasets like ImageNet, offering significant runtime advantages and enabling scalable, efficient deployment.
+(a) MNIST
+(b) ImageNet
+Figure 7: Runtime comparison of different data processing modes.
+D.2.2 Regression Task
+We evaluate the computational efficiency and predictive performance of TorchCP and an R-based package, AdaptiveConformal (Susmann et al., 2023) on a time series regression task, using the France electricity forecasting dataset [^13] https://cran.r-project.org/web/packages/opera/index.html.
+Specifically, we use the first 350 time steps to train a SARIMA model with order (1, 0, 0) and seasonal order (1, 1, 0), and construct prediction intervals for the remaining time steps.
+To assess computational efficiency, we apply the ACI algorithm(Gibbs and Candes, 2021) using both TorchCP and AdaptiveConformal.
+Both AdaptiveConformal and TorchCP are executed on CPU in this experiment to ensure a fair comparison.
+Moreover, since ACI performs recalibration for each test instance, the reported time includes both calibration and inference.
+Figure 8(a) reports the runtime results of these two packages.
+The results show that TorchCP achieves significantly higher computational efficiency compared to the R-based package, reducing running time from 6.1 seconds to around 0.06 seconds.
+Figures 8(b) and 8(c) illustrate the prediction intervals produced by AdaptiveConformal and TorchCP, respectively.
+The two sets of intervals are nearly identical, demonstrating the correctness and consistency of the TorchCP implementation.
+This substantial improvement highlights the efficiency advantage of TorchCP for time series regression.
+(a) Time
+(b) AdaptiveConformal
+(c) TorchCP
+Figure 8: Comparison of time consumption and prediction intervals between AdaptiveConformal and TorchCP on the France electricity dataset.
+Appendix E The Advantage of our modular design
+The modular design of TorchCP prioritizes low coupling to ensure flexibility in integrating diverse training algorithms, non-conformity score functions, and prediction workflows, accommodating the wide variety of CP methods and tasks.
+In contrast, previous packages (e.g., PUNCC and Fortuna) focus on a monolithic design where predictors and scoring functions are often predefined within a specific predictor (e.g., Split CP).
+This limits extensibility when new CP predictors or custom score functions are needed.
+TorchCP ’s decoupled design allows users to independently select or define new trainers, scores, or predictors, enabling flexible integration of different components.
+In what follows, we elaborate on the concrete benefits of our modular design:
+  • Flexible combination: users can freely combine Trainer, Score, and Predictor modules to customize conformal prediction workflows.
+For instance, a practitioner can use different non-conformity scores (e.g., LAC or APS) within any predictor or trainer, while other packages only implement scores for predictors.
+  • Modification isolation: our module design ensures that changes to one module (e.g., refining the APS score) do not affect other modules.
+This isolation minimizes the risk of unintended side effects, reducing maintenance complexity and preventing errors from propagating across the system.
+For instance, tightly coupled systems like PUNCC require redefining the predictor to modify a scoring function, while TorchCP confines changes to the target module, streamlining updates and preserving stability.
+  • Research extensibility: by adhering to standardized interfaces, new functionality (e.g., a novel non-conformity score) can be integrated into TorchCP without altering existing modules.
+This contrasts with PUNCC, where introducing new scoring methods demands extensive predictor modifications.
+Extensibility supports rapid iteration and long-term evolvability, enabling efficient adaptation to new research or applications.
+  • Ease of testing and debugging: decoupled modules can be tested and debugged independently, as their interactions are governed by well-defined interfaces.
+This modularity simplifies identifying and resolving issues within a single module (e.g., debugging a Trainer algorithm) without navigating complex dependencies.
+In TorchCP, this reduces testing overhead and accelerates development cycles compared to tightly coupled systems, where debugging often requires analyzing the entire pipeline.

### v4

@@
 Abstract
-Conformal prediction (CP) is a robust statistical framework that generates prediction intervals or sets with guaranteed coverage probability, addressing the challenge of quantifying predictive uncertainty in deep learning.
-Despite advancements in deep learning architectures and datasets, reliable uncertainty estimation remains elusive, making CP increasingly vital.
-This paper introduces TorchCP, a PyTorch-native library designed to integrate state-of-the-art CP algorithms into deep learning tasks, including classification, regression, graph neural networks, and large language models.
-TorchCP offers a comprehensive suite of advanced methodologies, a modular design for easy customization, and full GPU-accelerated scalability.
-Released under the LGPL-3.0 license, TorchCP has gained widespread adoption with over 12,582 PyPi downloads.
-It is supported by approximately 16,132 lines of code, 564 unit tests achieving 100% coverage, and comprehensive documentation.
-By bridging statistics and computer science, TorchCP empowers researchers and practitioners to advance conformal prediction in diverse deep learning applications.
+Conformal prediction (CP) is a powerful statistical framework that generates prediction intervals or sets with guaranteed coverage probability.
+While CP algorithms have evolved beyond traditional classifiers and regressors to sophisticated deep learning models like deep neural networks (DNNs), graph neural networks (GNNs), and large language models (LLMs), existing CP libraries often lack the model support and scalability for large-scale DL scenarios.
+This paper introduces TorchCP, a PyTorch-native library designed to integrate state-of-the-art CP algorithms into deep learning techniques, including DNN-based classifier/regressor, GNN, and LLM.
+Released under the LGPL-3.0 license, TorchCP comprises about 16k lines of code, validated with 100% unit test coverage and detailed documentation.
+Notably, TorchCP enables CP-specific training algorithms, online prediction, and GPU-accelerated batch processing, achieving up to 90% reduction in inference time on large datasets.
+With its low-coupling design, comprehensive suite of advanced methods, and full GPU scalability, TorchCP empowers researchers and practitioners to enhance uncertainty quantification across cutting-edge applications.
 Keywords: Conformal Prediction, Deep Learning
 1 Introduction
-Conformal prediction (CP) is a powerful statistical framework generating prediction intervals or sets that contain true labels with a specified coverage probability (Vovk et al., 2005).
-In recent years, deep learning has evolved dramatically with advanced model architectures and vast datasets, yet it struggles to reliably measure predictive uncertainty, making conformal prediction an increasingly attractive solution.
-Beyond classification and regression, conformal prediction methods have been tailored for diverse deep learning tasks, such as graph neural networks (GNNs) for molecular modeling and large language models (LLMs) for natural language understanding.
-This underscores the need for a comprehensive and efficient CP toolbox tailored for deep learning models, facilitating robust and accessible conformal prediction methods across diverse deep learning applications.
-This work presents TorchCP, a PyTorch-native library of conformal prediction for deep learning.
-Firstly, TorchCP implements a wide range of state-of-the-art CP algorithms encompassing both training and inference methods designed for deep learning tasks, e.g., classification (ordinal classification), regression (time series), GNN, and LLM.
-Secondly, the library’s modular design allows developers to easily use or customize each component (e.g., losses or scores) of the library.
-Lastly, TorchCP is fully built on PyTorch with GPU support and batch processing, enabling seamless scalability for large-scale deep learning tasks.
-The TorchCP library, released under the LGPL-3.0 license, can be installed via PyPi with the command ‘‘pip install torchcp’’ and has been downloaded 12,582 times  [^1] https://pepy.tech/projects/torchcp.
-The implementation consists of approximately 16,132 lines of code and is thoroughly tested with 564 unit tests, achieving 100% test coverage.
-We also provide comprehensive online documentation  [^2] https://torchcp.readthedocs.io/en/latest/ to guide users through TorchCP’s features.
-We hope the library can bridge the gap between statistics and computer science, enabling researchers and practitioners from both backgrounds to advance conformal prediction in deep learning.
-2 The library design and implementation
-Figure 1: An illustration of the modular design of TorchCP.
-2.1 Modular design
-For low coupling, we separate the TorchCP library into separate modules that can be easily maintained, tested, and integrated.
-For each task, the code implementation of the library contains 3 modules: Trainer, Score, and Predictor.
-Specifically, Trainer defines training algorithms for improving conformal prediction, along with loss functions.
-Score defines non-conformity score functions that quantify how different a new sample is from the training data, e.g., LAC (Sadinle et al., 2019).
-Predictor defines CP workflows (e.g., calibration and prediction), such as Split CP  [^3] A detailed introduction of Split CP is presented in Appendix A.1 (Papadopoulos et al., 2002).
-The modular design allows a straightforward extension to new tasks or new CP methods that arise in this active research field.
-We visualize the modular design of TorchCP in Figure 1, highlighting the relationships among the three modules.
-The Trainer module is used to train or fine-tune deep models, employed by the Predictor module for calibration and prediction.
-The Score module implements non-conformity score functions, enabling their flexible use across different trainers and predictors.
-This modular structure empowers users to independently select or customize trainers, score functions, and predictors for diverse conformal prediction tasks, significantly improving flexibility and maintainability.
-Additionally, TorchCP achieves strong extensibility by providing standardized module interfaces, making it straightforward to define new methods.
-Appendix B provides guidelines for adding customized modules using the provided interfaces, while Appendix E discusses the advantages of TorchCP ’s modular design.
-2.2 Implementation
-Table 1: The CP algorithms implemented in TorchCP.
-Task Component Methods
-Classification Trainer&Loss ConfTr (Stutz et al., 2022), ConfTS (Xi et al., 2025), C-Adapter (Liu et al., 2024), UA Loss (Einbinder et al., 2022), SCPO (Bellotti, 2021)
+Conformal prediction (CP) is a powerful statistical framework generating prediction intervals or sets with guaranteed coverage probability (Vovk et al., 2022), offering a distribution-free solution to quantify predictive uncertainty.
+While CP algorithms have been developed beyond traditional classifiers and regressors to sophisticated DL models like graph neural networks (GNNs) and large language models (LLMs), existing CP libraries often lack the model support and scalability for adapting to large-scale DL scenarios.
+This underscores the need for a comprehensive and scalable CP toolbox tailored for deep learning models, facilitating robust and accessible conformal prediction methods across cutting-edge applications.
+This work presents TorchCP, a PyTorch-native library that advances conformal prediction for deep learning by implementing a wide range of state-of-the-art CP algorithms, including training and inference methods tailored for models such as DNN classifier/regressor, GNN, and LLM.
+Its low-coupling design facilitates easy customization of components like trainers or scores, while full GPU support and batch processing ensure seamless scalability for large-scale tasks.
+TorchCP excels over existing libraries by delivering superior speed, with up to 90% faster inference times on large datasets compared to MAPIE and PUNCC (Appendix D.2), while also encompassing a broader algorithmic coverage than TorchUQ.
+This PyTorch-native integration and extensibility empower statisticians and DL practitioners to seamlessly apply CP across cutting-edge applications with enhanced DL compatibility.
+2 Main contributions
+Conformal prediction constructs prediction sets $C(X)$ for input $X$, ensuring the true outcome $Y$ satisfies $\mathbb{P}(Y\in\mathcal{C}(X))\geq 1-\alpha$, where $\alpha$ is a miscoverage rate.
+Existing CP libraries encounter significant challenges in deep learning, including limited native support for modern models (like DNN, GNN, and LLM), poor computational efficiency when scaling to large datasets, and monolithic designs that restrict the flexible customization for diverse applications.
+TorchCP overcomes these shortcomings through a PyTorch-native framework that improves model compatibility, efficiency, and adaptability, thereby enhancing CP’s applicability to DL.
+Its low-coupling design and GPU-accelerated batch processing provide a robust foundation for tackling complex, real-world problems, detailed in the following parts.
+Figure 1: The modules of TorchCP.
+2.1 Advantages for DL scenarios
+Low-coupling design Low-coupling design is one of the most fundamental concepts in object-oriented design (McNatt and Bieman, 2001), which minimizes the dependency a class has on other classes.
+While many libraries (e.g., PUNCC) are implemented in a high-coupling design, TorchCP is structured into three core modules: Trainer, Score, and Predictor, each tailored to specific aspects of CP workflows.
+The Trainer module implements CP-specific training algorithms, such as ConfTr (Stutz et al., 2022), which optimizes deep models to improve efficiency.
+The Score module calculates non-conformity measures, e.g., LAC (Sadinle et al., 2019), to quantify how atypical a new sample is relative to training data.
+The Predictor module implements CP workflows, including calibration and prediction, to generate outputs, e.g., prediction sets/intervals, p-values, and conformal predictive distributions.
+This low-coupling design is visualized in Figure 1, illustrating the interplay among modules.
+The Score module is used to create various score objects that serve as inputs for both the Trainer and Predictor modules.
+The Trainer module optimizes DL models using CP-specific loss functions.
+The Predictor module then integrates the trained model with the score objects to generate conformal outputs, such as prediction sets, intervals, p-values, or predictive distributions.
+In this manner, TorchCP achieves strong reusability and extensibility by providing standardized module interfaces, making it straightforward to define new methods.
+Appendix B provides guidelines for adding customized modules using the provided interfaces, while Appendix E discusses the advantages of TorchCP ’s low-coupling design.
+Table 1: The CP algorithms implemented in TorchCP for various deep learning models.
+Model types Component CP Algorithms
+DNN Classifier Trainer&Loss ConfTr (Stutz et al., 2022), ConfTS (Xi et al., 2025), C-Adapter (Liu et al., 2024), UA Loss (Einbinder et al., 2022), SCPO (Bellotti, 2021)
 Score Function LAC (Sadinle et al., 2019), APS (Romano et al., 2020), RAPS (Angelopoulos et al., 2020), SAPS (Huang et al., 2024), KNN (Gammerman and Vovk, 2007), Margin (Löfström et al., 2015), TOPK (Angelopoulos et al., 2020)
-Predictor Split CP (Papadopoulos et al., 2002) Class-conditional CP (Vovk, 2012), Clustered CP (Ding et al., 2024), Weighted CP (Tibshirani et al., 2019), RC3P (Shi et al., 2024)
-Regression Score Function ABS (Papadopoulos et al., 2002), NORABS (Vovk et al., 2005), CQR (Romano et al., 2019), CQRR (Sesia and Candès, 2020), CQRM (Sesia and Candès, 2020), CQRFM (Kivaranovic et al., 2020), R2CCP (Guha et al., 2023)
-Predictor Ensemble CP (Xu and Xie, 2021), ACI (Gibbs and Candes, 2021), AgACI (Zaffran et al., 2022)
+Predictor Split CP (Papadopoulos et al., 2002),Class-conditional CP (Vovk, 2012), Clustered CP (Ding et al., 2024), Weighted CP (Tibshirani et al., 2019), RC3P (Shi et al., 2024)
+DNN Regressor Score Function ABS (Papadopoulos et al., 2002), NORABS (Papadopoulos et al., 2011), CQR (Romano et al., 2019), CQRR (Sesia and Candès, 2020), CQRM (Sesia and Candès, 2020), CQRFM (Kivaranovic et al., 2020), R2CCP (Guha et al., 2023)
+Predictor Split CP, Ensemble CP (Xu and Xie, 2021), ACI (Gibbs and Candes, 2021), AgACI (Zaffran et al., 2022)
 GNN Trainer&Loss CF-GNN (Huang et al., 2023)
@@
 Methodological scope
-Table 1 outlines the CP algorithms included in TorchCP across 4 deep learning tasks: classification, regression (time series), GNN, and LLM , along with their corresponding evaluation metrics in Table 4.
-In total, TorchCP provides 6 training algorithms, 16 score functions, and 10 predictors tailored to various tasks.
+Table 1 outlines the CP algorithms in TorchCP across 4 types of DL models [^1] Model types are defined by their input/output data differences, impacting the choice of CP algorithms.: DNN classifier/regressor, GNN, and LLM, along with their evaluation metrics in Table 4.
+Current libraries are typically limited in algorithmic coverage, e.g., MAPIE only includes split CP.
+In TorchCP, we provide 6 training algorithms, 17 scores, and 10 predictors.
 A brief overview of all supported methods and their advantages is available in Appendix A.
@@
 TorchCP is a PyTorch-native library equipped with GPU acceleration and batch processing capabilities.
-It enables efficient computation on large datasets and scales seamlessly to real-world deep learning applications.
+It enables efficient computation on large datasets and scales seamlessly to deep learning applications, e.g., achieving up to 90% reduction in inference time on ImageNet (See Figure 6(b)).
 Appendix D.2 presents empirical evidence of its efficiency advantages on large-scale datasets.
-Tutorials
-TorchCP aims to provide a user-friendly framework for conformal prediction in deep learning.
-The library contains illustrative code showcasing functionality for common tasks (See the “examples” folder  [^4] https://github.com/ml-stat-Sustech/TorchCP/tree/master/examples), and Appendix C offers a practical guide to help users identify relevant examples.
-In Appendix D, we provide reproduced results of classical CP methods, confirming the correctness of our implementation.
-Below, we present a generic example  [^5] The full code of the example is available at https://github.com/ml-stat-Sustech/TorchCP/blob/master/examples/classification_conftr_cifar100.py to show the usability of TorchCP.
-First, we employ ConfTr (Stutz et al., 2022) to train a model for conformal prediction (optional, using standard trained models also work):
+2.2 Tutorials
+TorchCP provides a user-friendly framework for conformal prediction in deep learning with a detailed documentation [^2] https://torchcp.readthedocs.io/en/latest/index.html.
+The library contains illustrative code showcasing functionality for common applications (See the “examples” folder  [^3] https://github.com/ml-stat-Sustech/TorchCP/tree/master/examples) with a practical guide in Appendix C.
+Moreover, in Appendix D, we provide reproduced results of classical CP methods, confirming the correctness of our implementation.
+Below, we present a generic example  [^4] The full source code is available at TorchCP CIFAR-100 Example.
+to show the usability of TorchCP.
+First, we utilize ConfTr (Stutz et al., 2022) to train a neural network, i.e., init_model, on a training dataset encapsulated in train_loader:
 1from torchcp.classification.trainer import ConfTrTrainer
-2trainer = ConfTrTrainer(init_model, alpha=0.1) # Create a trainer with a NN
+2trainer = ConfTrTrainer(init_model, alpha=0.1) # Initialize trainer with model
 3trained_model = trainer.train(train_loader, num_epochs=10) # Training 10 epochs
 With the trained model, users can proceed with conformal prediction by selecting an appropriate score function and predictor.
-As an example, we implement Split CP with a miscoverage rate of $\alpha=0.1$ using the LAC score :
+Below, we implement Split CP with a miscoverage rate of $\alpha=0.1$, using the LAC score on calibration data encapsulated in cal_loader:
 1predictor = SplitPredictor(score_function=LAC(), model=trained_model, alpha=0.1)
-2predictor.calibrate(cal_loader) # Compute the threshold on calibrate data
-3predict_set = predictor.predict(test_instance) # Predict the set for an instance
+2predictor.calibrate(cal_loader) # Set threshold with calibration data
+3predict_set = predictor.predict(test_instance) # Predict set for test instance
 3 Related library
-Conformal Prediction (CP) is a distribution-free framework for quantifying the uncertainty of machine learning model predictions.
-While some uncertainty quantification libraries, such as PUNCC (Mendil et al., 2023), implement certain CP algorithms, their limited efficiency, scalability, and algorithmic coverage restrict their use in deep learning.
-For instance, most libraries support only a narrow range of tasks, such as regression and classification, and lack support for complex applications like generative modeling or time-series forecasting.
-Additionally, the absence of GPU acceleration and batch processing in these libraries makes them inefficient for large-scale scenarios.
-In contrast, TorchCP offers a clean interface, enhanced extensibility, high efficiency, and a comprehensive set of CP algorithms.
-For example, TorchCP reduces inference time by up to 90% compared to PUNCC on large-scale datasets (see Appendix D.2 for details).
-We summarize the comparison of different libraries in Table 2 and discuss the advantages of TorchCP ’s modular design in Appendix E.
-Table 2: Feature comparison of open-source libraries supporting conformal prediction. “Training” refers to CP-specific training algorithms, and “Online” denotes online CP.
-Library Task Support Functionality Support Backend
-Classification Regression GNN LLM Training Online GPU Batch
-nonconformist (Linusson et al., 2020) Numpy
-TorchUQ (Zhao et al., 2024) PyTorch
-Crepes (Boström, 2022) Numpy
-PUNCC (Mendil et al., 2023) Numpy
-MAPIE (Cordier et al., 2023) Numpy
-Fortuna (Detommaso et al., 2024) JAX
-TorchCP (ours) PyTorch
+Conformal prediction is a distribution-free framework for quantifying uncertainty, yet existing libraries face notable limitations in deep learning applications.
+Libraries such as nonconformist, PUNCC, and MAPIE offer restricted native support for modern models like GNNs and LLMs, relying on NumPy without GPU acceleration, which hinders efficient scaling to large datasets.
+Moreover, tools like TorchUQ and Fortuna provide some GPU support but fall short in algorithmic coverage, e.g., TorchUQ supports only regressors.
+Additionally, many libraries like PUNCC are implemented in a high-coupling design, where predictors and score functions are tightly coupled within a single class.
+TorchCP addresses these gaps with its PyTorch-native integration, supporting a broader range of modern models and offering a modular, extensible architecture, as detailed in Table 8.
+We present a detailed comparison of the relevant libraries in Appendix F.
+Table 2: Comparison of Open-Source CP Libraries  [^7] The definition of model support and functionality support is presented in Appendix F..
+“Classifier” and “Regressor” refer to DNN-based classifiers and regressors [^8] TorchCP also supports traditional machine learning models like random forests but is primarily designed for advanced deep learning models within the PyTorch ecosystem..
+The term “Training” denotes support for CP-specific training, while “Online” refers to capabilities for semi-online conformal prediction.
+“$p$-value” and “CPD” denote outputting $p$-values and conformal predictive distributions, respectively.
+Library Model Support Functionality Support Coupling Backend
+Classifier Regressor GNN LLM Training Online GPU Batch $p$-value CPD
+nonconformist (Linusson et al., 2020) low Numpy
+TorchUQ (Zhao et al., 2024) low PyTorch
+Crepes (Boström, 2022) low Numpy
+PUNCC (Mendil et al., 2023) high Numpy
+MAPIE (Cordier et al., 2023) high Numpy
+Fortuna (Detommaso et al., 2024) high JAX
+TorchCP (ours) low PyTorch
 4 Conclusion and Future Work
-In this work, we introduced TorchCP, a PyTorch-native library that seamlessly integrates conformal prediction (CP) into deep learning workflows.
-Supporting a wide range of state-of-the-art CP algorithms, TorchCP leverages PyTorch’s GPU acceleration and batch processing, offering a modular design for easy customization and user-friendly documentation for accessibility.
+In this work, we introduced TorchCP, a PyTorch-native library that seamlessly integrates conformal prediction algorithms into deep learning workflows.
+Supporting a wide range of state-of-the-art CP algorithms, TorchCP leverages PyTorch’s GPU acceleration and batch processing, offering a low-coupling design for easy customization and user-friendly documentation for accessibility.
 Looking ahead, we plan to expand TorchCP by incorporating emerging CP algorithms for evolving architectures like diffusion models and multimodal models, while optimizing for real-time performance.
@@
 Component Method Advantage/Use case
-Classification
+DNN Classifier
 Score LAC To achieve the smallest prediction sets
@@
 ConfTS Improving efficiency via temperature scaling
-Regression
+DNN Regressor
 Score ABS Simple and interpretable residual-based scores
-NORABS Improving robustness of prediction intervals via noise-aware normalization
+NORABS Enhancing Prediction Interval Robustness through Normalized Scoring
 CQR Flexible interval width via quantile regression
@@
 A.1 Split CP
-Conformal prediction (Papadopoulos et al., 2002; Balasubramanian et al., 2014; Manokhin, 2022; Angelopoulos et al., 2023, 2024) is a statistical framework that generates uncertainty sets/intervals containing ground-truth labels with a desired probability.
-Split Conformal Prediction (Papadopoulos et al., 2002) (Split CP) is the most widely-used version of the conformal prediction procedure.
+Conformal prediction (Papadopoulos et al., 2002; Balasubramanian et al., 2014; Manokhin, 2022; Angelopoulos et al., 2023, 2024) is a statistical framework that generates prediction sets/intervals containing ground-truth labels with a desired probability.
+Split Conformal Prediction (Vovk et al., 2022) (Split CP) is the most widely-used version of the conformal prediction procedure.
 Specifically, Split CP divides a given dataset into two disjoint subsets: one for training the base model and the other for conformal calibration.
@@
     $\mathcal{C}(\bm{x}_{n+1})=\{y:V(\bm{x}_{n+1},y)\leq\widehat{Q}_{1-\alpha}\}.$
-A.2 Classification
-Here, we provide a brief overview of conformal prediction methods for multi-class classification tasks as supported by TorchCP.
+A.2 DNN Classifier
+Here, we provide a brief overview of conformal prediction methods implemented using a DNN classifier with the TorchCP library.
 Let $\mathcal{X}\subset\mathbb{R}^{d}$ be the input space and $\mathcal{Y}:=\{1,\dots,K\}$ be the label space.
 We use $(X,Y)\sim\mathcal{P}_{\mathcal{X}\mathcal{Y}}$ to denote a random data pair satisfying a joint data distribution $\mathcal{P}_{\mathcal{X}\mathcal{Y}}$ and $f:\mathcal{X}\rightarrow\mathbb{R}^{K}$ to denote a classification neural network.
-Thus, the classifier $\pi:\mathcal{X}\rightarrow\Delta^{K-1}$is defined as $\sigma\circ f$, where $\Delta^{K-1}$ is a ($K$-1)-dimensional probability simplex and $\sigma$ is a normalization function such as the softmax function.
+Thus, the classifier $\pi:\mathcal{X}\rightarrow\Delta^{K-1}$is defined as $\sigma\circ f$, where $\Delta^{K-1}$ is a (KK-1)-dimensional probability simplex and $\sigma$ is a normalization function such as the softmax function.
 Let $\pi_{y}(x)$ be the probability of $y$ conditional on $x$, i.e., $P(y|x)$.
@@
 RAPS (Angelopoulos et al., 2020): RAPS adds a regularization to penalize noisy tail probabilities and regularize the number of samples in the uncertainty set:
-$V(\bm{x},y,u;\pi):=\sum\limits_{y^{\prime}\in\mathcal{Y}}\pi_{y^{\prime}}(\bm{x})\mathds{1}(r_{f}(\bm{x},y^{\prime})<r_{f}(\bm{x},y))+u\cdot\pi_{y}(\bm{x})+\eta\cdot(r_{f}(\bm{x},y)-k_{reg})^{+},$
-where $\eta$ represents the weight of regularization, $k_{reg}\geq 0$ are regularization hyper-parameters and $(z)^{+}$ denotes the positive part of $z$.
+$V(\bm{x},y,u;\pi):=\sum\limits_{y^{\prime}\in\mathcal{Y}}\pi_{y^{\prime}}(\bm{x})\mathds{1}(r_{f}(\bm{x},y^{\prime})<r_{f}(\bm{x},y))+u\cdot\pi_{y}(\bm{x})+\lambda\cdot(r_{f}(\bm{x},y)-k)^{+},$
+where $\lambda$ represents the weight of regularization, $k\geq 0$ are regularization hyper-parameters and $(z)^{+}$ denotes the positive part of $z$.
 TOPK (Angelopoulos et al., 2020): TOPK is a score function based solely on the ranking information: $V(\bm{x},y,u):=\sum\limits_{y^{\prime}\in\mathcal{Y}}\mathds{1}(r_{f}(\bm{x},y^{\prime})<r_{f}(\bm{x},y))+u$.
@@
 where $V(\bm{x},y,u):=\sum\limits_{y^{\prime}\in\mathcal{Y}}\pi_{y^{\prime}}(\bm{x})\mathds{1}(r_{f}(\bm{x},y^{\prime})<r_{f}(\bm{x},y))+u\cdot\pi_{y}(\bm{x})$.
-A.3 Regression
-In the context of regression, the goal is to predict a continuous target variable $y\in\mathbb{R}$ based on an input $\bm{x}\in\mathcal{X}\subset\mathbb{R}^{d}$, where $(X,Y)\sim\mathcal{P}_{\mathcal{X}\mathcal{Y}}$ follows a joint data distribution $\mathcal{P}_{\mathcal{X}\mathcal{Y}}$.
+A.3 DNN Regressor
+For DNN Regressor, the goal is to predict a continuous target variable $y\in\mathbb{R}$ based on an input $\bm{x}\in\mathcal{X}\subset\mathbb{R}^{d}$, where $(X,Y)\sim\mathcal{P}_{\mathcal{X}\mathcal{Y}}$ follows a joint data distribution $\mathcal{P}_{\mathcal{X}\mathcal{Y}}$.
 Let $f:\mathcal{X}\to\mathcal{F}$ denote a regression model trained on a dataset $\{(\bm{x_{i}},y_{i})\}_{i=1}^{n}$, where $\mathcal{F}$ represents the space of model outputs, $\bm{x}_{i}$ are the features and $y_{i}$ are the observed target values.
@@
 Then the prediction set for a new instance $\bm{x}_{n+1}$ is constructed by $\mathcal{C}(\bm{x}_{n+1})=[f(\bm{x}_{n+1})-\widehat{Q}_{1-\alpha},f(\bm{x}_{n+1})+\widehat{Q}_{1-\alpha}].$
-NORABS (Vovk et al., 2005): The NORABS score $V(\bm{x},y):=\frac{|f(\bm{x})-y|}{\hat{\sigma}(\bm{x})}$ is a normalized variant of ABS.
-It uses models $f(\bm{x})$ and $\hat{\sigma}(\bm{x})$ trained to maximize the likelihood of the data with respect to $\mathbb{E}[Y_{\text{test}}|X_{\text{test}}=\bm{x}]$ and $\sqrt{\text{Var}[Y_{\text{test}}|X_{\text{test}}=\bm{x}]}$, respectively.
-Then the prediction set for a new instance $\bm{x}_{n+1}$ is constructed by $\mathcal{C}(\bm{x}_{n+1})=[f(\bm{x}_{n+1})-\widehat{Q}_{1-\alpha}\cdot\hat{\sigma}(\bm{x}_{n+1}),f(\bm{x}_{n+1})+\widehat{Q}_{1-\alpha}\cdot\hat{\sigma}(\bm{x}_{n+1})].$
-CQR (Romano et al., 2019): The CQR score $V(\bm{x},y):=\max(f(\bm{x})_{\alpha/2}-y,\;y-$$f(\bm{x})_{1-\alpha/2})$ measures the maximum absolute deviation between the true value $y$ and the predicted quantile interval $[f(\bm{x})_{\alpha/2},f(\bm{x})_{1-\alpha/2}]$.
+NORABS (Papadopoulos et al., 2011; Papadopoulos and Haralambous, 2011): The NORABS score $V(\bm{x},y):=\frac{|f(\bm{x})-y|}{\gamma+\xi(\bm{x})}$ normalizes the regression error using difficulty measure deviation $\xi(\bm{x})$.
+The prediction set for $\bm{x}_{n+1}$ is $\mathcal{C}(\bm{x}_{n+1})=[f(\bm{x}_{n+1})-\widehat{Q}_{1-\alpha}\cdot(\gamma+\xi(\bm{x}_{n+1})),f(\bm{x}{n+1})+\widehat{Q}_{1-\alpha}\cdot(\gamma+\xi(\bm{x}_{n+1}))]$.
+Additionally, we provide the normalized score for Gaussian Process Regression (Papadopoulos, 2024).
+CQR (Romano et al., 2019): The CQR score $V(\bm{x},y):=\max(f(\bm{x})_{\alpha/2}-y,\;y$-$f(\bm{x})_{1-\alpha/2})$ measures the maximum absolute deviation between the true value $y$ and the predicted quantile interval $f(\bm{x})_{\alpha/2},f(\bm{x})_{1-\alpha/2}]$.
 Then the prediction set for a new instance $\bm{x}_{n+1}$ is constructed by: $\mathcal{C}(\bm{x}_{n+1})=[f(\bm{x})_{\alpha/2}-\widehat{Q}_{1-\alpha},\;f(\bm{x})_{1-\alpha/2}+\widehat{Q}_{1-\alpha}].$
 CQRR (Sesia and Candès, 2020): The CQRR score extends the standard CQR method by introducing a scaling mechanism.
-This score normalizes the deviations between the true value $y$ and the predicted quantile interval $[f(\bm{x})_{\alpha/2},f(\bm{x})_{1-\alpha/2}]$ by the width of the interval itself.
+This score normalizes the deviations between the true value $y$ and the predicted quantile interval $f(\bm{x})_{\alpha/2},f(\bm{x})_{1-\alpha/2}]$ by the width of the interval itself.
 Formally, for a data pair $(\bm{x},y)$, the CQRR score is defined as $V(\bm{x},y):=\max\left(\frac{f(\bm{x})_{\alpha/2}-y}{\hat{s}_{\alpha}(\bm{x})},\;\frac{y-f(\bm{x})_{1-\alpha/2}}{\hat{s}_{\alpha}(\bm{x})}\right),$ where $\hat{s}_{\alpha}(\bm{x}):=f(\bm{x})_{1-\alpha/2}-f(\bm{x})_{\alpha/2}+\epsilon$ is the scaling factor, and $\epsilon>0$ is a small constant added to prevent division by zero.
@@
 CQRM (Sesia and Candès, 2020): The CQRM score extends the CQRR score by introducing two scaling factors to independently normalize the deviations for the lower and upper bounds of the prediction interval.
-This method requires the model to output predictions at three quantile levels: $[\alpha/2,1/2,1-\alpha/2]$.
+This method requires the model to output predictions at three quantile levels: $\alpha/2,1/2,1-\alpha/2]$.
 The two scaling factors are defined as $\hat{s}_{\alpha,\text{lo}}(\bm{x}):=f(\bm{x})_{1/2}-f(\bm{x})_{\alpha/2}+\epsilon$ and $\hat{s}_{\alpha,\text{up}}(\bm{x}):=f(\bm{x})_{1-\alpha/2}-f(\bm{x})_{1/2}+\epsilon$.
@@
 $\mathcal{C}(\bm{x}_{n+1})=\left[f(\bm{x}_{n+1})_{\alpha/2}-\widehat{Q}_{1-\alpha}\cdot\hat{s}_{\alpha,\text{lo}}(\bm{x}_{n+1}),\;f(\bm{x}_{n+1})_{1-\alpha/2}+\widehat{Q}_{1-\alpha}\cdot\hat{s}_{\alpha,\text{up}}(\bm{x}_{n+1})\right].$
-CQRFM (Kivaranovic et al., 2020): The CQRFM score is similar to CQRM, as it also requires the model to output predicted quantiles at $[\alpha/2,1/2,1-\alpha/2]$.
+CQRFM (Kivaranovic et al., 2020): The CQRFM score is similar to CQRM, as it also requires the model to output predicted quantiles at $\alpha/2,1/2,1-\alpha/2]$.
 However, it normalizes the deviations from the median $f(\bm{x})_{1/2}$ by the fractional distance to the nearest quantile bounds.
@@
 A.4 GNN
-We consider the prediction tasks on graph data.
+In this section, we discuss the applications of conformal prediction algorithms on graph data.
 The Graph is defined as $\mathcal{G}=(\mathcal{V},\mathcal{E})$, where $\mathcal{V}:=\{v_{i}\}_{i=1}^{N}$ denotes the node set and $\mathcal{E}$ denotes the edge set with $|\mathcal{E}|=E$.
@@
 The label of $v_{i}$ is $y_{i}\in\mathcal{Y}$, where $\mathcal{Y}:=\{1,2,...,K\}$ denotes the label space.
-DAPS (Zargarbashi et al., 2023): DAPS proposes a diffusion-based method that incorporates neighborhood information by leveraging the network homophily.
-Specifically, the score function is defined as: $V(\bm{x}_{i},y,\lambda):=(1-\lambda)\hat{V}(\bm{x}_{i},y)+\frac{\lambda}{|\mathcal{N}_{i}|}\sum_{v_{j}\in\mathcal{N}_{i}}\hat{V}(\bm{x}_{j},y),$ where $\hat{V}(\cdot,\cdot)$ is the basic non-conformity score function, e.g., APS, and $\lambda$ is a diffusion parameter.
+[…]
 SNAPS (Song et al., 2024): SNAPS, which is essentially an enhanced version of the DAPS, aggregates non-conformity scores of nodes with high feature similarity to ego node and one-hop graph structural neighbors.
@@
 A.6 Metrics
-In this section, we provide a brief overview of the evaluation metrics supported by TorchCP for various predictive tasks, as summarized in Table 4.
+In this section, we provide a brief overview of the evaluation metrics supported by TorchCP for diverse modern models, as summarized in Table 4.
 These metrics facilitate rigorous and fine-grained assessment of prediction sets, offering deep insights into model uncertainty.
 Table 4: Implemented evaluation metrics in TorchCP.
-Task Metrics
-Classification Coverage, Size, CovGap (Ding et al., 2024), WSC (Romano et al., 2020), SSCV (Angelopoulos et al., 2020), Violated classes (Kasa and Taylor, 2023), DiffViolation (Angelopoulos et al., 2020), , etc.
-Regression Coverage, Size
+Deep Models Metrics
+DNN Classifier Coverage, Size, CovGap (Ding et al., 2024), WSC (Romano et al., 2020), SSCV (Angelopoulos et al., 2020), Violated classes (Kasa and Taylor, 2023), DiffViolation (Angelopoulos et al., 2020), S Criterion, N Criterion, U Criterion, F Criterion, M Criterion, F Criterion, OU Criterion, OF Criterion, OM Criterion, OE Criterion (Vovk et al., 2022), etc.
+DNN Regressor Coverage, Size
 GNN Coverage, Size, Singleton ratio (Zargarbashi et al., 2023)
@@
   1. 1.
-    Classification.
+    DNN classifier.
 This example code [^10] https://github.com/ml-stat-Sustech/TorchCP/blob/master/examples/classification_splitcp_cifar100.py applies Split CP with the LAC score on CIFAR-100 using a pretrained ResNet-20.
@@
   2. 2.
-    Regression.
-The example code of regression [^8] https://github.com/ml-stat-Sustech/TorchCP/blob/master/examples/regression_cqr_synthetic.py shows the complete workflow of applying Split CP with the Conformalized Quantile Regression score (Romano et al., 2019) on a synthetic dataset in regression task, including model training, calibration, prediction interval generation, and evaluation.
+    DNN Regressor.
+The example code of regression [^11] https://github.com/ml-stat-Sustech/TorchCP/blob/master/examples/regression_cqr_synthetic.py shows the complete workflow of applying Split CP with the Conformalized Quantile Regression score (Romano et al., 2019) on a synthetic dataset in a regression task, including model training, calibration, prediction interval generation, and evaluation.
   3. 3.
     GNN.
-The node classification example [^9] https://github.com/ml-stat-Sustech/TorchCP/blob/master/examples/gnn_transductive_coraml.py implements a Graph Convolutional Network (GCN) (Kipf and Welling, 2017) for node classification on the CoraML dataset (McCallum et al., 2000) under a transductive setting.
+The GNN example [^12] https://github.com/ml-stat-Sustech/TorchCP/blob/master/examples/gnn_transductive_coraml.py implements a Graph Convolutional Network (GCN) (Kipf and Welling, 2017) for node classification on the CoraML dataset (McCallum et al., 2000) under a transductive setting.
 Conformal prediction is then applied to generate prediction sets with guaranteed coverage.
@@
 The LLM example [^13] https://github.com/ml-stat-Sustech/TorchCP/blob/master/examples/llm_ConformalLM_TriviaQA.py demonstrates how to deploy conformal prediction in large language model generation tasks, i.e., TriviaQA (Joshi et al., 2017).
-C.2 Example codes for online conformal prediction
-Online conformal prediction means that after predicting a test point, its true label is immediately revealed, and the test sample is incorporated into the calibration set.
+C.2 Example codes for Semi-online conformal prediction
+Semi-online conformal prediction means that after predicting a test point, its true label is immediately revealed, and the test sample is incorporated into the calibration set.
 TorchCP supports this setting as well, with the following examples:
   1. 1.
-    Regression.
-In TorchCP, the online setup of the regression task is integrated into ACIPredictor (Gibbs and Candes, 2021).
+    DNN regressor.
+In TorchCP, the semi-online setup of the regression task is integrated into ACIPredictor (Gibbs and Candes, 2021).
 Therefore, we added an example code  [^14] https://github.com/ml-stat-Sustech/TorchCP/blob/master/examples/timeseries_aci_synthetic.py for the regression task in TorchCP.
   2. 2.
-    Classification.
-For classification tasks, this online calibration approach can be naturally implemented in TorchCP by iteratively invoking the “calibrate()” function with new incoming data.
-To clarify this, we have added an online conformal prediction example 121212 https://github.com/ml-stat-Sustech/TorchCP/blob/master/examples/classification_splitcp_cifar100_online.py for a classification task in TorchCP.
+    DNN Classifier.
+For DNN classification tasks, this semi-online calibration approach can be naturally implemented in TorchCP by iteratively invoking the “calibrate()” function with new incoming data.
+To clarify this, we have added a semi-online conformal prediction example 151515 https://github.com/ml-stat-Sustech/TorchCP/blob/master/examples/classification_splitcp_cifar100_online.py for a classification task in TorchCP.
 Appendix D Empirical Results
@@
 D.1 Evaluation results
-Classification.
+DNN Classifier.
 In Figure 2, we present the experimental results of Split CP on ImageNet, using LAC, APS, RAPS (penalty=0.1, kreg =0), and SAPS (weight=0.2).
@@
 Figure 3: Conditional coverage results on ImageNet image classification with $\alpha=0.1$.
-Regression.
+DNN regressor.
 In Figure 4, we present an example of experimental results on the Community dataset (Redmond, 2002), using ABS, CQR, CQRR, CQRM, CQRFM, R2CCP with Split CP. The number of trials is 1000.
@@
 Figure 7: Runtime comparison of different data processing modes.
+(a) Time
+(b) AdaptiveConformal
+(c) TorchCP
+Figure 8: Comparison of time consumption and prediction intervals between AdaptiveConformal and TorchCP on the France electricity dataset.
 D.2.2 Regression Task
@@
 This substantial improvement highlights the efficiency advantage of TorchCP for time series regression.
-(a) Time
-(b) AdaptiveConformal
-(c) TorchCP
-Figure 8: Comparison of time consumption and prediction intervals between AdaptiveConformal and TorchCP on the France electricity dataset.
-Appendix E The Advantage of our modular design
-The modular design of TorchCP prioritizes low coupling to ensure flexibility in integrating diverse training algorithms, non-conformity score functions, and prediction workflows, accommodating the wide variety of CP methods and tasks.
-In contrast, previous packages (e.g., PUNCC and Fortuna) focus on a monolithic design where predictors and scoring functions are often predefined within a specific predictor (e.g., Split CP).
+Appendix E The Advantage of our low-coupling design
+Low-coupling design is one of the most fundamental concepts in object-oriented design, which minimizes the dependency a class has on other classes.
+The low-coupling design of TorchCP ensures flexibility in integrating diverse training algorithms, non-conformity score functions, and prediction workflows, accommodating the wide variety of CP methods and tasks.
+In contrast, most current packages focus on a high-coupling design where predictors and scoring functions are often predefined within a specific predictor (e.g., Split CP).
 This limits extensibility when new CP predictors or custom score functions are needed.
 TorchCP ’s decoupled design allows users to independently select or define new trainers, scores, or predictors, enabling flexible integration of different components.
-In what follows, we elaborate on the concrete benefits of our modular design:
+In what follows, we elaborate on the concrete benefits of our low-coupling design:
   • Flexible combination: users can freely combine Trainer, Score, and Predictor modules to customize conformal prediction workflows.
 For instance, a practitioner can use different non-conformity scores (e.g., LAC or APS) within any predictor or trainer, while other packages only implement scores for predictors.
-  • Modification isolation: our module design ensures that changes to one module (e.g., refining the APS score) do not affect other modules.
+Moreover, some libraries (e.g., TorchUQ) implemented various non-conformity scores as functions, but this style suffers from extending to those scores with extra hyperparameters (e.g., the $k$ and $\lambda$ of RAPS).
+  • Modification isolation: our low-coupling design ensures that changes to one module (e.g., refining the APS score) do not affect other modules.
 This isolation minimizes the risk of unintended side effects, reducing maintenance complexity and preventing errors from propagating across the system.
@@
   • Research extensibility: by adhering to standardized interfaces, new functionality (e.g., a novel non-conformity score) can be integrated into TorchCP without altering existing modules.
-This contrasts with PUNCC, where introducing new scoring methods demands extensive predictor modifications.
+Differently, MAPIE follows a high-coupling design, making it challenging to introduce new predictors, like class-conditional CP.
 Extensibility supports rapid iteration and long-term evolvability, enabling efficient adaptation to new research or applications.
@@
 In TorchCP, this reduces testing overhead and accelerates development cycles compared to tightly coupled systems, where debugging often requires analyzing the entire pipeline.
+Appendix F Comparison with related libraries
+Numerous libraries implement conformal prediction (CP) algorithms, yet their suitability for deep learning (DL) scenarios differs significantly, particularly in supporting modern DL models like GNNs and LLMs, as well as their scalability for large-scale datasets.
+In Table 8, we present a concise comparison of open-source CP libraries, focusing on supported model types and functionalities:
+  • Model type refers to the types of machine learning models the library can work with.
+They are classified based on differences in input and output data, which influence the selection of applicable CP algorithms.
+  • Functionality describes the operational capabilities of the library, such as how it processes data, computational optimization, or prediction forms.
+In the following, we compare several popular CP libraries, emphasizing TorchCP’s superior compatibility and performance with DL applications.
+nonconformist (Linusson et al., 2020): A NumPy-based library supporting basic CP algorithms for classification and regression.
+It lacks GPU acceleration, batch processing, and compatibility with DL models like DNN Classifier, DNN Regressor, GNNs, or LLMs, making it unsuitable for complex DL scenarios.
+While nonconformist implements predictors and scores as different classes following a low-coupling design, its algorithmic coverage is limited, implementing only inductive CP and aggregated CP with 5 scores.
+TorchUQ (Zhao et al., 2024): A PyTorch library focused on uncertainty quantification, including some CP methods for regression, with GPU support.
+However, it lacks comprehensive algorithm implementation, omitting CP algorithms for classifiers and CP-specific training, and does not support advanced deep learning models like GNNs or LLMs, limiting its applicability for complex deep learning Scenarios.
+TorchUQ uses a dictionary to map names to score functions in predictors, enabling developers to define new scores without modifying predictor logic.
+However, this style struggles to accommodate scores with additional hyperparameters (e.g., $k$ and $\lambda$ in RAPS), requiring extra logic to manage these variations, which increases coupling and complexity.
+Crepes (Boström, 2022): a NumPy-based library with extensive prediction forms that supports conformal classifiers, regressors, and predictive systems with semi-online CP, excelling in statistical classification and regression tasks.
+However, it lacks GPU acceleration, batch processing, and seamless extensions to modern models like GNNs or LLMs, limiting scalability for engineers tackling large-scale or complex deep learning problems.
+Similar to TorchUQ, Crepes implements non-conformity scores as functions passed to predictors, so it also suffers from supporting scores with additional hyperparameters (e.g., $k$ and $\lambda$ in RAPS).
+PUNCC (Mendil et al., 2023): A scikit-learn-compatible library delivers Conformal Prediction for tasks like regression, classification, and anomaly detection, but struggles with modern DL applications.
+It does not support advanced DL models such as DNN, GNNs, or LLMs, lacks GPU acceleration and batch processing, and performs significantly slower on large datasets (up to 10x slower than TorchCP in ImageNet, see Figure 6(b)).
+PUNCC defines nonconformity scores as functions, but ties each score to a dedicated predictor, resulting in high coupling.
+MAPIE (Cordier et al., 2023): A scikit-learn-based library providing CP sets/intervals for classical classification/regression tasks.
+It excels in traditional machine learning workflows but lacks support for deep learning models, such as DNNs, GNNs, or LLMs.
+It also does not offer online CP, GPU acceleration, and batch processing.
+MAPIE implements scores as different classes with a predict function, making it challenging to support other predictors (like class-conditional CP (Vovk, 2012)).
+Fortuna (Detommaso et al., 2024): a JAX-based library, supports CP methods in deep learning with GPU acceleration and select conformal prediction methods.
+It enhances scalability for deep learning but has limited algorithmic coverage, lacking training algorithms and online CP, and does not support GNNs or LLMs.
+Additionally, JAX’s steeper learning curve may challenge statisticians familiar with PyTorch.
+Like PUNCC, Fortuna defines a dedicated classifier for each score, leading to high-coupling design.
+In contrast, TorchCP, a PyTorch-native library, significantly enhances conformal prediction for deep learning by providing a comprehensive suite of state-of-the-art CP algorithms.
+It offers exceptional extensibility through low-coupling design with three modules (See Appendix E) and unified interfaces across various algorithms (See Appendix B).
+Notably, TorchCP seamlessly integrates CP algorithms with advanced deep learning models, such as GNNs and LLMs, and leverages GPU-accelerated batch processing to ensure scalability for large-scale deep learning scenarios (See Appendix D.2).

### v5

@@
 Conformal prediction (CP) is a powerful statistical framework that generates prediction intervals or sets with guaranteed coverage probability.
-While CP algorithms have evolved beyond traditional classifiers and regressors to sophisticated deep learning models like deep neural networks (DNNs), graph neural networks (GNNs), and large language models (LLMs), existing CP libraries often lack the model support and scalability for large-scale DL scenarios.
-This paper introduces TorchCP, a PyTorch-native library designed to integrate state-of-the-art CP algorithms into deep learning techniques, including DNN-based classifier/regressor, GNN, and LLM.
+While CP algorithms have evolved beyond traditional classifiers and regressors to sophisticated deep learning models like deep neural networks (DNNs), graph neural networks (GNNs), and large language models (LLMs), existing CP libraries often lack the model support and scalability for large-scale deep learning (DL) scenarios.
+This paper introduces TorchCP, a PyTorch-native library designed to integrate state-of-the-art CP algorithms into DL techniques, including DNN-based classifiers/regressors, GNNs, and LLMs.
 Released under the LGPL-3.0 license, TorchCP comprises about 16k lines of code, validated with 100% unit test coverage and detailed documentation.
@@
 1 Introduction
-Conformal prediction (CP) is a powerful statistical framework generating prediction intervals or sets with guaranteed coverage probability (Vovk et al., 2022), offering a distribution-free solution to quantify predictive uncertainty.
-While CP algorithms have been developed beyond traditional classifiers and regressors to sophisticated DL models like graph neural networks (GNNs) and large language models (LLMs), existing CP libraries often lack the model support and scalability for adapting to large-scale DL scenarios.
-This underscores the need for a comprehensive and scalable CP toolbox tailored for deep learning models, facilitating robust and accessible conformal prediction methods across cutting-edge applications.
-This work presents TorchCP, a PyTorch-native library that advances conformal prediction for deep learning by implementing a wide range of state-of-the-art CP algorithms, including training and inference methods tailored for models such as DNN classifier/regressor, GNN, and LLM.
+Conformal prediction (CP) is a statistical framework generating prediction intervals or sets with guaranteed coverage probability (Vovk et al., 2022), offering a distribution-free solution to quantify predictive uncertainty.
+While CP algorithms have been developed beyond traditional classifiers and regressors to sophisticated deep learning (DL) models like graph neural networks (GNNs) and large language models (LLMs), existing CP libraries often lack the model support and scalability required for large-scale DL scenarios.
+This underscores the need for a comprehensive and scalable CP toolbox tailored for DL models, facilitating robust and accessible conformal prediction methods across cutting-edge applications.
+This work presents TorchCP, a PyTorch-native library that advances conformal prediction for DL by implementing a wide range of state-of-the-art CP algorithms, including training and inference methods tailored for models such as DNN classifiers/regressors, GNNs, and LLMs.
 Its low-coupling design facilitates easy customization of components like trainers or scores, while full GPU support and batch processing ensure seamless scalability for large-scale tasks.
-TorchCP excels over existing libraries by delivering superior speed, with up to 90% faster inference times on large datasets compared to MAPIE and PUNCC (Appendix D.2), while also encompassing a broader algorithmic coverage than TorchUQ.
+TorchCP excels over existing libraries by delivering superior speed, with up to 90% faster inference times on large datasets compared to MAPIE and PUNCC (Appendix F), while also encompassing a broader algorithmic coverage than TorchUQ.
 This PyTorch-native integration and extensibility empower statisticians and DL practitioners to seamlessly apply CP across cutting-edge applications with enhanced DL compatibility.
 2 Main contributions
-Conformal prediction constructs prediction sets $C(X)$ for input $X$, ensuring the true outcome $Y$ satisfies $\mathbb{P}(Y\in\mathcal{C}(X))\geq 1-\alpha$, where $\alpha$ is a miscoverage rate.
-Existing CP libraries encounter significant challenges in deep learning, including limited native support for modern models (like DNN, GNN, and LLM), poor computational efficiency when scaling to large datasets, and monolithic designs that restrict the flexible customization for diverse applications.
+Conformal prediction constructs a prediction set $C(X)$ for an input $X$, ensuring the true outcome $Y$ satisfies $P(Y\in\mathcal{C}(X))\geq 1-\alpha$, where $\alpha$ is a miscoverage rate.
+Existing CP libraries encounter significant challenges in DL, including limited native support for modern models (like DNNs, GNNs, and LLMs), poor computational efficiency when scaling to large datasets, and monolithic designs that restrict the flexible customization for diverse applications.
 TorchCP overcomes these shortcomings through a PyTorch-native framework that improves model compatibility, efficiency, and adaptability, thereby enhancing CP’s applicability to DL.
@@
 The Trainer module implements CP-specific training algorithms, such as ConfTr (Stutz et al., 2022), which optimizes deep models to improve efficiency.
-The Score module calculates non-conformity measures, e.g., LAC (Sadinle et al., 2019), to quantify how atypical a new sample is relative to training data.
+The Score module calculates non-conformity measures, e.g., LAC (Sadinle et al., 2019), to quantify how atypical a new sample is relative to the training data.
 The Predictor module implements CP workflows, including calibration and prediction, to generate outputs, e.g., prediction sets/intervals, p-values, and conformal predictive distributions.
@@
 The Trainer module optimizes DL models using CP-specific loss functions.
-The Predictor module then integrates the trained model with the score objects to generate conformal outputs, such as prediction sets, intervals, p-values, or predictive distributions.
+The Predictor module then integrates the trained model with the score objects to generate conformal outputs, such as prediction sets/intervals, p-values, or predictive distributions.
 In this manner, TorchCP achieves strong reusability and extensibility by providing standardized module interfaces, making it straightforward to define new methods.
 Appendix B provides guidelines for adding customized modules using the provided interfaces, while Appendix E discusses the advantages of TorchCP ’s low-coupling design.
-Table 1: The CP algorithms implemented in TorchCP for various deep learning models.
 Model types Component CP Algorithms
-DNN Classifier Trainer&Loss ConfTr (Stutz et al., 2022), ConfTS (Xi et al., 2025), C-Adapter (Liu et al., 2024), UA Loss (Einbinder et al., 2022), SCPO (Bellotti, 2021)
+DNN Classifier Trainer&Loss ConfTr (Stutz et al., 2022), ConfTS (Xi et al., 2025), C-Adapter (Liu et al., 2025), UA Loss (Einbinder et al., 2022), SCPO (Bellotti, 2021)
 Score Function LAC (Sadinle et al., 2019), APS (Romano et al., 2020), RAPS (Angelopoulos et al., 2020), SAPS (Huang et al., 2024), KNN (Gammerman and Vovk, 2007), Margin (Löfström et al., 2015), TOPK (Angelopoulos et al., 2020)
@@
 LLM Predictor Conformal LLM (Quach et al., 2024)
+Table 1: The CP algorithms implemented in TorchCP for various deep learning models.
 Methodological scope
-Table 1 outlines the CP algorithms in TorchCP across 4 types of DL models [^1] Model types are defined by their input/output data differences, impacting the choice of CP algorithms.: DNN classifier/regressor, GNN, and LLM, along with their evaluation metrics in Table 4.
+Table 1 outlines the CP algorithms in TorchCP across 4 types of DL models [^1] Model types are defined by their input/output data differences, impacting the choice of CP algorithms.: DNN classifiers/regressors, GNNs, and LLMs, along with evaluation metrics in Table 4.
 Current libraries are typically limited in algorithmic coverage, e.g., MAPIE only includes split CP.
@@
 TorchCP is a PyTorch-native library equipped with GPU acceleration and batch processing capabilities.
-It enables efficient computation on large datasets and scales seamlessly to deep learning applications, e.g., achieving up to 90% reduction in inference time on ImageNet (See Figure 6(b)).
+It enables efficient computation on large datasets and scales seamlessly to DL applications, e.g., achieving up to 90% reduction in inference time on ImageNet (See Figure 6(b)).
 Appendix D.2 presents empirical evidence of its efficiency advantages on large-scale datasets.
@@
 3 Related library
-Conformal prediction is a distribution-free framework for quantifying uncertainty, yet existing libraries face notable limitations in deep learning applications.
-Libraries such as nonconformist, PUNCC, and MAPIE offer restricted native support for modern models like GNNs and LLMs, relying on NumPy without GPU acceleration, which hinders efficient scaling to large datasets.
+Conformal prediction is a distribution-free framework for quantifying predictive uncertainty, yet existing libraries face notable limitations in DL applications.
+Libraries, such as nonconformist, PUNCC, and MAPIE, offer restricted native support for modern models like GNNs and LLMs, relying on NumPy without GPU acceleration, which hinders efficient scaling to large datasets.
 Moreover, tools like TorchUQ and Fortuna provide some GPU support but fall short in algorithmic coverage, e.g., TorchUQ supports only regressors.
@@
 We present a detailed comparison of the relevant libraries in Appendix F.
-Table 2: Comparison of Open-Source CP Libraries  [^7] The definition of model support and functionality support is presented in Appendix F..
-“Classifier” and “Regressor” refer to DNN-based classifiers and regressors [^8] TorchCP also supports traditional machine learning models like random forests but is primarily designed for advanced deep learning models within the PyTorch ecosystem..
-The term “Training” denotes support for CP-specific training, while “Online” refers to capabilities for semi-online conformal prediction.
-“$p$-value” and “CPD” denote outputting $p$-values and conformal predictive distributions, respectively.
 Library Model Support Functionality Support Coupling Backend
@@
 TorchCP (ours) low PyTorch
+Table 2: Comparison of Open-Source CP Libraries  [^7] The definitions of model support and functionality support are presented in Appendix F..
+“Classifier” and “Regressor” refer to DNN-based classifiers and regressors [^8] TorchCP also supports traditional machine learning models like random forests, but is primarily designed for advanced DL models within the PyTorch ecosystem..
+The term “Training” denotes support for CP-specific training, while “Online” refers to capabilities for semi-online conformal prediction.
+“$p$-value” and “CPD” denote outputting $p$-values and conformal predictive distributions, respectively.
 4 Conclusion and Future Work
@@
 Through open-source collaboration on GitHub, we aim to foster community-driven innovation, empowering reliable AI in fields such as healthcare, finance, and autonomous systems.
+Acknowledgments and Disclosure of Funding
+Hongxin Wei is supported by the Shenzhen Fundamental Research Program (Grant No.
+JCYJ20230807091809020).
+Bingyi Jing’s research is partly supported by NSFC 12371290.
+We gratefully acknowledge the support of the Center for Computational Science and Engineering at the Southern University of Science and Technology for our research.
 Appendix A Details of Conformal Prediction Methods in TorchCP
@@
 Furthermore, Table 3 offers practical guidelines to help users select appropriate methods based on their specific application needs.
-Table 3: Advantages or use cases of CP methods across different tasks.
 Component Method Advantage/Use case
@@
 KNN A score function based solely on feature-space nearest neighbors without a predictive model.
-Predictor Class-conditional CP To achieve coverage for every class
+Predictor Class-conditional CP To achieve coverage for each class
 Clustered CP For class-conditional coverage with limited data per class
@@
 Predictor Conformal LLM To achieve statistically guaranteed prediction sets for language models through calibrated sampling
+Table 3: Advantages or use cases of CP methods across different tasks.
 A.1 Split CP
@@
   4. 4.
-    Use the conformal threshold $\widehat{Q}_{1-\alpha}$ to generate a prediction set for a new instance $\bm{x}_{n+1}$:
+    Use the threshold $\widehat{Q}_{1-\alpha}$ to generate a prediction set for a new instance $\bm{x}_{n+1}$:
     $\mathcal{C}(\bm{x}_{n+1})=\{y:V(\bm{x}_{n+1},y)\leq\widehat{Q}_{1-\alpha}\}.$
@@
 $\mathcal{L}_{\mathrm{ConfTS}}=(\widehat{Q}_{1-\alpha}-V(\bm{x},y))^{2}$
-C-Adapter (Liu et al., 2024): C-Adapter proposes to optimize an adapter with a Conformal Discriminative (CD) Loss.
+C-Adapter (Liu et al., 2025): C-Adapter proposes to optimize an adapter with a Conformal Discriminative (CD) Loss.
 The CD loss is constructed to improve the $\alpha$-free efficiency of prediction sets: $\mathcal{L}_{\mathrm{CD}}=\frac{1}{K}\sum_{j\in\mathcal{Y}}\{\sigma(V(\bm{x},y)-V(\bm{x},j))\}$.
@@
 NORABS (Papadopoulos et al., 2011; Papadopoulos and Haralambous, 2011): The NORABS score $V(\bm{x},y):=\frac{|f(\bm{x})-y|}{\gamma+\xi(\bm{x})}$ normalizes the regression error using difficulty measure deviation $\xi(\bm{x})$.
-The prediction set for $\bm{x}_{n+1}$ is $\mathcal{C}(\bm{x}_{n+1})=[f(\bm{x}_{n+1})-\widehat{Q}_{1-\alpha}\cdot(\gamma+\xi(\bm{x}_{n+1})),f(\bm{x}{n+1})+\widehat{Q}_{1-\alpha}\cdot(\gamma+\xi(\bm{x}_{n+1}))]$.
+The prediction set for $\bm{x}_{n+1}$ is $\mathcal{C}(\bm{x}_{n+1})=[f(\bm{x}_{n+1})-\widehat{Q}_{1-\alpha}\cdot(\gamma+\xi(\bm{x}_{n+1})),f(\bm{x}_{n+1})+\widehat{Q}_{1-\alpha}\cdot(\gamma+\xi(\bm{x}_{n+1}))]$.
 Additionally, we provide the normalized score for Gaussian Process Regression (Papadopoulos, 2024).
@@
 Based on these scores, R2CCP outputs the interval where all non-conformity scores are smaller than $\widehat{Q}_{1-\alpha}$.
-Ensemble CP (Xu and Xie, 2021): Ensemble CP trains multiple models based on randomly sampled subsets of the data and aggregates their nonconformity scores using a specific aggregation function, such as mean or median.
+Ensemble CP (Xu and Xie, 2021): Ensemble CP trains multiple models based on randomly sampled subsets of the data and aggregates their non-conformity scores using a specific aggregation function, such as mean or median.
 It can update $\widehat{Q}_{1-\alpha}$ over time, making it adaptable to dynamic time series.
@@
 These metrics facilitate rigorous and fine-grained assessment of prediction sets, offering deep insights into model uncertainty.
-Table 4: Implemented evaluation metrics in TorchCP.
 Deep Models Metrics
@@
 LLM SetLoss, SSCL (Quach et al., 2024)
+Table 4: Implemented evaluation metrics in TorchCP.
 Appendix B Extensibility
@@
 Appendix C Example Codes for TorchCP
-TorchCP provides a collection of example codes in the “examples" folder  [^9] https://github.com/ml-stat-Sustech/TorchCP/tree/master/examples of the official codebase, designed to help users quickly get started with various conformal prediction tasks.
+TorchCP provides a collection of example codes in the “examples” folder  [^9] https://github.com/ml-stat-Sustech/TorchCP/tree/master/examples of the official codebase, designed to help users quickly get started with various conformal prediction tasks.
 In this section, we briefly introduce representative examples for the four supported tasks.
@@
 For DNN classification tasks, this semi-online calibration approach can be naturally implemented in TorchCP by iteratively invoking the “calibrate()” function with new incoming data.
-To clarify this, we have added a semi-online conformal prediction example 151515 https://github.com/ml-stat-Sustech/TorchCP/blob/master/examples/classification_splitcp_cifar100_online.py for a classification task in TorchCP.
+To clarify this, we have added a semi-online conformal prediction example  [^15] https://github.com/ml-stat-Sustech/TorchCP/blob/master/examples/classification_splitcp_cifar100_online.py for a classification task in TorchCP.
 Appendix D Empirical Results
@@
 DNN Classifier.
-In Figure 2, we present the experimental results of Split CP on ImageNet, using LAC, APS, RAPS (penalty=0.1, kreg =0), and SAPS (weight=0.2).
-In Figure 3, we present a comparative analysis of different conformal prediction methods on the ImageNet dataset, including Split CP, Class-conditional CP, and Clustered CP with APS.
-The number of trials is 1000.
+In Figure 2, we present the experimental results of Split CP on ImageNet, using LAC, APS, RAPS ($\lambda$=0.1, $k$ =0), and SAPS ($\lambda$=0.2).
+Figure 3 compares Split CP, Class-conditional CP, and Clustered CP on the ImageNet dataset, each employing the APS scoring rule.
+Results are averaged over 1,000 independent trials.
 (a) Coverage
@@
 The generation process terminates after the fifth sample, as the cumulative score surpasses the conformal threshold.
-Table 5: An Example of prediction sets from Conformal LLM on TriviaQA.
 Question Men Against the Sea and Pitcairn’s Island were two sequels to what famous novel?
@@
 Answer 5 Robinson Crusoe 0.1384 False
+Table 5: An Example of prediction sets from Conformal LLM on TriviaQA.
 D.2 Comparison of computational efficiency
@@
 We compared the computational efficiency of five libraries: PUNCC, MAPIE, Fortuna, Crepes, and TorchCP (GPU-based full-batch processing).
-In the dataset, we use 48,000 samples for training, 6,000 for calibration, and 6,000 for testing.
-For TorchCP, we use a LeNet-5 model for training and inference using PyTorch on a GPU.
+For the dataset, we use 48,000 samples for training, 6,000 for calibration, and 6,000 for testing.
+For TorchCP, we use a LeNet-5 model for training and inference using PyTorch on the GPU.
 For PUNCC, MAPIE, and Crepes, we obtain model predictions using PyTorch on a GPU, and then convert output probabilities to NumPy arrays for conformal calibration and inference using CPU.
@@
 Figure 6: Runtime comparison of different conformal prediction libraries.
-Choices of Batch processing.
-To assess the efficiency of two processing modes (Full-batch and batch), we conduct experiments with TorchCP on both MNIST and ImageNet, using the same data splits and models as in prior experiments.
-We utilize GPU-accelerated computing for all experiments except the full-batch mode on ImageNet, where the dataset’s size exceeds GPU memory capacity.
-Figures 7(a) and  7(b) present the running times of different data processing modes on MNIST and ImageNet, respectively.
-On the small-scale dataset - MNIST, using full-batch processing on GPU achieves the fastest implementation among the three choices, requiring only 26% computational time of that with CPU.
-When it comes to the ImageNet dataset, GPU-based batch processing completes calibration and inference in approximately 104.29 seconds, compared to 1022.55 seconds for CPU-based processing, achieving a 9.8× speedup.
-These findings suggest that GPU-based full-batch processing is optimal for small-scale datasets where GPU memory is sufficient to load the whole dataset.
-Conversely, GPU-based batch processing is preferable for large-scale datasets like ImageNet, offering significant runtime advantages and enabling scalable, efficient deployment.
+[…]
 (a) MNIST
@@
 D.2.2 Regression Task
-We evaluate the computational efficiency and predictive performance of TorchCP and an R-based package, AdaptiveConformal (Susmann et al., 2023) on a time series regression task, using the France electricity forecasting dataset [^16] https://cran.r-project.org/web/packages/opera/index.html.
+We evaluate the computational efficiency and predictive performance of TorchCP and an R-based CP package, AdaptiveConformal (Susmann et al., 2023) on a time series regression task, using the France electricity forecasting dataset [^16] https://cran.r-project.org/web/packages/opera/index.html.
 Specifically, we use the first 350 time steps to train a SARIMA model with order (1, 0, 0) and seasonal order (1, 1, 0), and construct prediction intervals for the remaining time steps.
@@
 In contrast, most current packages focus on a high-coupling design where predictors and scoring functions are often predefined within a specific predictor (e.g., Split CP).
-This limits extensibility when new CP predictors or custom score functions are needed.
+This limits extensibility when new CP predictors or user-defined score functions are needed.
 TorchCP ’s decoupled design allows users to independently select or define new trainers, scores, or predictors, enabling flexible integration of different components.
@@
   • Flexible combination: users can freely combine Trainer, Score, and Predictor modules to customize conformal prediction workflows.
-For instance, a practitioner can use different non-conformity scores (e.g., LAC or APS) within any predictor or trainer, while other packages only implement scores for predictors.
+For instance, a practitioner can use different non-conformity scores (e.g., LAC or APS) within any predictor or trainer, while other packages only implement specified scores for predictors.
 Moreover, some libraries (e.g., TorchUQ) implemented various non-conformity scores as functions, but this style suffers from extending to those scores with extra hyperparameters (e.g., the $k$ and $\lambda$ of RAPS).
@@
 Appendix F Comparison with related libraries
-Numerous libraries implement conformal prediction (CP) algorithms, yet their suitability for deep learning (DL) scenarios differs significantly, particularly in supporting modern DL models like GNNs and LLMs, as well as their scalability for large-scale datasets.
+Numerous libraries implement conformal prediction (CP) algorithms, yet their suitability for deep learning scenarios differs significantly, particularly in supporting modern deep learning models like GNNs and LLMs, as well as their scalability for large-scale datasets.
 In Table 8, we present a concise comparison of open-source CP libraries, focusing on supported model types and functionalities:
@@
   • Functionality describes the operational capabilities of the library, such as how it processes data, computational optimization, or prediction forms.
-In the following, we compare several popular CP libraries, emphasizing TorchCP’s superior compatibility and performance with DL applications.
-nonconformist (Linusson et al., 2020): A NumPy-based library supporting basic CP algorithms for classification and regression.
-It lacks GPU acceleration, batch processing, and compatibility with DL models like DNN Classifier, DNN Regressor, GNNs, or LLMs, making it unsuitable for complex DL scenarios.
+In the following, we compare several popular CP libraries, emphasizing TorchCP’s superior compatibility and performance with deep learning applications.
+nonconformist (Linusson et al., 2020): A NumPy-based library supporting basic CP algorithms for classification and regression tasks.
+However, it lacks GPU acceleration, batch processing, and compatibility with deep learning models like DNN Classifiers, DNN Regressors, GNNs, or LLMs, making it unsuitable for complex deep learning scenarios.
 While nonconformist implements predictors and scores as different classes following a low-coupling design, its algorithmic coverage is limited, implementing only inductive CP and aggregated CP with 5 scores.
@@
 However, this style struggles to accommodate scores with additional hyperparameters (e.g., $k$ and $\lambda$ in RAPS), requiring extra logic to manage these variations, which increases coupling and complexity.
-Crepes (Boström, 2022): a NumPy-based library with extensive prediction forms that supports conformal classifiers, regressors, and predictive systems with semi-online CP, excelling in statistical classification and regression tasks.
+Crepes (Boström, 2022): a NumPy-based library with extensive prediction forms that supports the classifiers, regressors, and predictive systems with semi-online CP, excelling in statistical classification and regression tasks.
 However, it lacks GPU acceleration, batch processing, and seamless extensions to modern models like GNNs or LLMs, limiting scalability for engineers tackling large-scale or complex deep learning problems.
 Similar to TorchUQ, Crepes implements non-conformity scores as functions passed to predictors, so it also suffers from supporting scores with additional hyperparameters (e.g., $k$ and $\lambda$ in RAPS).
-PUNCC (Mendil et al., 2023): A scikit-learn-compatible library delivers Conformal Prediction for tasks like regression, classification, and anomaly detection, but struggles with modern DL applications.
-It does not support advanced DL models such as DNN, GNNs, or LLMs, lacks GPU acceleration and batch processing, and performs significantly slower on large datasets (up to 10x slower than TorchCP in ImageNet, see Figure 6(b)).
-PUNCC defines nonconformity scores as functions, but ties each score to a dedicated predictor, resulting in high coupling.
+PUNCC (Mendil et al., 2023): A scikit-learn-compatible library delivers Conformal Prediction for tasks like regression, classification, and anomaly detection, but struggles with modern deep learning applications.
+It does not support advanced deep learning models such as DNNs, GNNs, or LLMs, lacks GPU acceleration and batch processing, and performs significantly slower on large datasets (up to 10x slower than TorchCP in ImageNet, see Figure 6(b)).
+PUNCC defines non-conformity scores as functions, but ties each score to a dedicated predictor, resulting in high coupling.
 MAPIE (Cordier et al., 2023): A scikit-learn-based library providing CP sets/intervals for classical classification/regression tasks.